### v1

# Panacea [Uncaptioned image]: Mitigating Harmful Fine-tuning for Large Language Models via Post-fine-tuning Perturbation

Abstract
Harmful fine-tuning attack introduces significant security risks to the fine-tuning services. Main-stream defenses aim to vaccinate the model such that the later harmful fine-tuning attack is less effective. However, our evaluation results show that such defenses are fragile– with a few fine-tuning steps, the model still can learn the harmful knowledge. To this end, we do further experiment and find that an embarrassingly simple solution– adding purely random perturbations to the fine-tuned model, can recover the model from harmful behaviors, though it leads to a degradation in the model’s fine-tuning performance. To address the degradation of fine-tuning performance, we further propose Panacea, which optimizes an adaptive perturbation that will be applied to the model after fine-tuning. Panacea maintains model’s safety alignment performance without compromising downstream fine-tuning performance. Comprehensive experiments are conducted on different harmful ratios, fine-tuning tasks and mainstream LLMs, where the average harmful scores are reduced by up-to 21.5%, while maintaining fine-tuning performance. As a by-product, we analyze the adaptive perturbation and show that different layers in various LLMs have distinct safety coefficients. Source code available at https://github.com/w-yibo/Panacea.
Machine Learning, ICML
1 Tsinghua University; 2 Shenzhen Campus of Sun Yat-sen University; 3 Didichuxing Co. Ltd; 4 Nanyang Technological University. Correspondence to: Li Shen <mathshenli@gmail.com>.
1 Introduction
Fine-tuning-as-a-service (OpenAI, ) is a popular business service to enhance model’s performance for customized datasets, domain-specific tasks, and private needs. However, recent studies (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023; Yi et al., 2024a; Li et al., 2024a; Ye et al., 2024) identify a safety issue, the harmful fine-tuning attack (Figure 1), where the model’s safety alignment is compromised when the fine-tuning dataset contains harmful data, even a small amount of harmful data can introduce significant security vulnerabilities. Moreover, harmful fine-tuning is often unintentional, as datasets may contain latent unsafe data that is difficult for users to detect. Since service providers are responsible for the harmful outputs generated by the model, there is a clear need for more effective solutions.
Figure 1: The harmful fine-tuning attack for fine-tuning-as-a-service scenarios. In the alignment stage, pretrained LLMs are aligned using alignment data to produce aligned LLMs. In the fine-tuning stage, aligned LLMs are further fine-tuned using fine-tuning data, which may contain harmful data, potentially leading to unsafe fine-tuned models.
Vaccine (Huang et al., 2024e) and RepNoise (Rosati et al., 2024a) are two representative defenses against the harmful fine-tuning attack. Vaccine improves the aligned model’s resistance towards harmful knowledge learning by solving a mini-max problem, while RepNoise aims to erase the harmful knowledge of the aligned model by perturbing the hidden representation given harmful data. However, our evaluation results show that with more fine-tuning steps, the vaccinated model produced by the two methods are still suffering from the negative effect of harmful fine-tuning attack – their harmful score are significantly increased, though with a slower rate compared to the baseline without defense.
Based on the above finding, it seems that the learning of harmful knowledge cannot be sufficiently suppressed before fine-tuning. From another angle, it may be worthwhile to consider a mitigation approach to the problem after fine-tuning. We start our exploration by a rather naive defense– adding purely random post-fine-tuning perturbation to the fine-tuned model. Our evaluation results surprisingly demonstrate that random perturbation can recover the model from harmful behavior, showing that such a naive method could be a potential defense. However, our subsequent evaluation shows that this method significantly degrades the fine-tuning performance of the model, indicating that such a method cannot strike a good balance between recovering from harmful behavior and maintaining fine-tuning performance. To this end, a subsequent question is that:
How to add post-fine-tuning perturbation to the fine-tuned model, such that it can be recovered from harmful behavior without hurting downstream task performance much?
Driven by this question, we propose Panacea (Figure 2), an iterative gradient-based method to iterately search for the post-fine-tuning perturbation. Panacea aims to solve a max-maximize optimization problem, such that the added perturbation can maximally increase the model’s harmful loss, ensuring that it can effectively recover the model from harmful behaviors. The experiment results show that for different harmful ratios during fine-tuning, our method’s average harmful score is reduced by 10.4% up to 21.5%, while fine-tuning performance improves by 0.3% over the standard alignment method. The ablation study on the adaptive perturbation show that it can reduce harmful scores by up to 23.2%, while maintaining competitive fine-tuning performance. The visualization experiments also reveal different layers in various LLMs have distinct safety coefficients, consistent with previous findings and providing additional evidence for layer-wise safety research.
The main contributions of this work contain three parts:
1. We find that adding purely random perturbations to the fine-tuned models could recover the model from harmful behavior, but does cause a loss of fine-tuning performance.
2. To mitigate harmful fine-tuning while maintaining the fine-tuning performance, we propose Panacea, a post-fine-tuning solution that formulates a max-max optimization problem, where the optimized perturbation maximally increases the harmful loss.
3. Experiments evaluate Panacea across diverse settings, demonstrating its effectiveness and generalization, while visualizations reveal safety coefficients across LLM layers.
2 Related Work
Safety Alignment. The safety alignment requires the model to output content that is both helpful and harmless, and to be able to output a refusal answer when given harmful instructions. Existing methods typically rely on supervised fine-tuning (SFT), RLHF (Ouyang et al., 2022), and variations of RLHF (e.g., PPO, DPO) (Bai et al., 2022; Wu et al., 2023; Dong et al., 2023; Rafailov et al., 2023; Yuan et al., 2023; Huang et al., 2021). These methods construct a safety-aligned dataset, and recent approaches focus on enhancing and better utilizing the aligned dataset (Liu et al., 2023a, b; Ye et al., 2023; Tekin et al., 2024).
Figure 2: Post-fine-tuning perturbation. The fine-tuned model exhibits a high harmful score (HS:$\downarrow$). Adding random perturbation reduces the harmful score but also decreases fine-tuning accuracy (FA:$\uparrow$). In contrast, incorporating our post-fine-tuning perturbation (see Algorithm 1) effectively lowers the harmful score while maintaining fine-tuning performance.
Harmful Fine-tuning. Fine-tuning-as-a-service becomes a mainstream method for LLMs API providers. Recent studies (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023; Li et al., 2024a; Ye et al., 2024; Lermen et al., 2023; He et al., 2024; Halawi et al., 2024; Chen et al., 2024; Hawkins et al., 2024; Poppi et al., 2024) show that LLMs trained with safety alignment can be jail-broken when fine-tuned on a dataset with a small amount of harmful data. In such cases, the model fails to refuse harmful instructions and outputs harmful responses. Many works (Rosati et al., 2024c; Leong et al., 2024; Anonymous, 2025c; Guo et al., 2024; Qi et al., 2024b; Rosati et al., 2024a) focus on analyzing the mechanisms of different harmful fine-tuning attacks. (Peng et al., 2024) proposes new safety metrics to evaluate harmful fine-tuning risk and (Rosati et al., 2024a) explores the safety risks when learning with reinforcement learning. Existing defenses can be categorized into three main categories (Huang et al., 2024c), i) Alignment stage solutions (Huang et al., 2024e; Rosati et al., 2024b; Tamirisa et al., 2024; Liu et al., 2024c; Pan et al., ; Huang et al., 2024b; Anonymous, 2025a; Liu et al., 2024a), ii) Fine-tuning stage solutions (Mukhoti et al., 2023; Bianchi et al., 2023; Zong et al., 2024; Wei et al., 2024; Huang et al., 2024d; Wang et al., 2024; Lyu et al., 2024; Eiras et al., 2024; Zhang et al., 2024; Li et al., 2024b; Shen et al., 2024; Anonymous, 2025b; Li et al., 2025; Du et al., 2024a; Choi et al., 2024; Luo et al., 2024; Qi et al., 2024a), iii) Post-fine-tuning stage solutions (Casper et al., 2024; Yi et al., 2024c; Du et al., 2024b; Hsu et al., 2024; Huang et al., 2024a; Zhu et al., 2024; Liu et al., 2024b; Wu et al., 2024; Yi et al., 2024b). The proposed method in this paper is applied in the post-fine-tuning stage, aiming to restore safety alignment without sacrificing fine-tuning performance.
Mainstream defenses focusing on the alignment stage lack sufficient durability against harmful fine-tuning (Qi et al., 2024b), motivating exploration of the post-fine-tuning stage. Existing post-fine-tuning solutions typically add perturbations based on prior knowledge, such as a safety subspace (Hsu et al., 2024) or safety-critical parameters (Zhu et al., 2024; Yi et al., 2024b). In contrast, Panacea optimizes an adaptive perturbation during fine-tuning without relying on prior knowledge.
3 Revisiting Alignment Defenses
3.1 Problem Setup
Harmful Fine-tuning. Harmful fine-tuning poses a significant threat to LLMs service providers (OpenAI, ). The scenario is illustrated in Figure 1, where LLMs service providers use an alignment dataset to perform safety alignment on a pretrained model, transforming it into an aligned model. Users then upload a fine-tuning dataset containing harmful data to the service provider. The fine-tuned dataset is deployed on the service provider’s server and used to generate personalized outputs for the users.
Threat Models. Following  (Qi et al., 2023; Rosati et al., 2024b; Huang et al., 2024b), we assume that, during the fine-tuning stage, $p$ (percentage) of the fine-tuning dataset consists of harmful data (i.e., harmful instruction-harmful response pairs), while the $1-p$ of data represents benign fine-tuning data (e.g., math question-answer pairs (Cobbe et al., 2021)). Furthermore, we assume that harmful and benign data are inseparable within the fine-tuning dataset.
Defense Assumptions. Assume that LLM providers host an alignment dataset (harmful instruction-harmless response pairs) used during the alignment stage. Such a dataset is also assumed to be available by Vaccine (Huang et al., 2024e), RepNoise (Rosati et al., 2024b), BEA(Wang et al., 2024). Additionally, we assume availability of a harmful dataset (harmful instruction-harmful response pairs). This harmful dataset is also assumed to be available by existing methods, e.g., RepNoise (Rosati et al., 2024b), TAR (Tamirisa et al., 2024), Booster (Huang et al., 2024b). Both the alignment dataset and the harmful dataset can be obtained from existing open-sourced datasets (e.g., BeaverTails).
3.2 Revisiting Alignment Defenses
In this section, we explore the existing alignment stage solutions during the harmful fine-tuning and discuss the feasibility of introducing random perturbation in the post-fine-tuning stage, thus inspiring our proposed method.
Figure 3: Model statistics (Left: harmful loss of three methods, Right: harmful score of three methods) after fine-tuning on fine-tuning dataset (10% of data is harmful) for different steps.
We select the SFT method (vanilla supervised fine-tuning with the alignment dataset), Vaccine (Huang et al., 2024e), and RepNoise (Rosati et al., 2024b), and perform fine-tuning on a fine-tuning dataset containing a small proportion (0.1) of harmful samples. As shown in Figure 3, the three methods start with different levels of harmful loss; however, as training progresses, they all achieve lower harmful loss, which corresponds to an increase in harmful score, making the model harmful. This naturally leads us to consider:
Can simply adding a random perturbation after the fine-tuning to increase the harmful loss to restore the safety alignment?
Figure 4: Model statistics (Left: harmful loss, Right: harmful $\downarrow$ and fine-tuning $\uparrow$) for fine-tuned model with noise intensities of 0 (no noise), 0.001, 0.01, 0.05, 0.1. (HS of 0.1 is greater than 80, FA of 0.1 is 0.7, and is not shown.)
To evaluate it, we design the experiments that add Gaussian noise with intensities of 0.001, 0.01, 0.05, and 0.1 to the fine-tuned model. The experimental results shown in Figure 4 indicate that adding random Gaussian noise increases harmful loss, demonstrating that random perturbations have the potential to prevent the harmful fine-tuning. We further measure the effects of adding random perturbations. As shown in Figure 4, the quantitative results reveal that random perturbations reduce the model’s harmful score. And the reduction effect improves as the noise intensity increases. However, excessive noise (0.1 of intensity) ultimately degrades the model. At the same time, random perturbations significantly impair the fine-tuned model’s performance.
Derived Insight. As the harmful dataset has the potential to direct the perturbation towards altering the harmful loss, thereby reducing perturbation’s impact on fine-tuning performance, we could optimize an perturbation with harmful dataset during fine-tuning stage, which recover the safety alignment while maintaining the fine-tuning performance.
4 Methodology
At the fine-tuning stage, we aim to optimize an adaptive perturbation that maximally increases the harmful loss. This perturbation is then added to the fine-tuned model after the fine-tuning process. Therefore, our method is a post-fine-tuning stage solution, formulated as a max-maximize optimization, as follows:
$\max_{\boldsymbol{w}}\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}\lambda(h(\boldsymbol{w}+\boldsymbol{\varepsilon})-h(\boldsymbol{w}))-g(\boldsymbol{w})$ (1)
where $\boldsymbol{w}$ and $\boldsymbol{\varepsilon}$ are the vanilla supervised aligned model parameters and the adaptive perturbation, respectively, $g(\boldsymbol{w})$ is the empirical loss over the fine-tuning dataset (contains harmful data) and $h(\boldsymbol{w})$ is the empirical loss over the harmful dataset. The outer optimization is maximizing the increase in harmful loss when adding the perturbation, while minimizing its fine-tuning loss. The term $h(\boldsymbol{w}+\boldsymbol{\varepsilon})-h(\boldsymbol{w})$ represents the harmful loss ascent when adding the perturbation, and $\lambda$ is a balancing hyper-parameter. The inner optimization $\max_{\boldsymbol{\varepsilon}}$ finds the optimal perturbation $\boldsymbol{\varepsilon}$ that maximizes the increase in harmful loss $h(\boldsymbol{w}+\boldsymbol{\varepsilon})$. The constraint $\|\boldsymbol{\varepsilon}\|\leq\rho$ ensures that the perturbation remains within a norm-bound $\rho$, preventing excessive perturbation.
To solve this max-maximize optimization problem, we adopt the alternative optimization. We alternatively solve the inner problem fixing $\boldsymbol{w}$ and solve the outer problem fixing $\boldsymbol{\varepsilon}$.
Close-form solution for the inner problem. Fixing $\boldsymbol{w}$, the inner optimization over $\boldsymbol{\varepsilon}$ could be solved with the following equation (See Appendix A for a proof):
$\boldsymbol{\varepsilon}_{t}^{*}=\rho\frac{\nabla h(\boldsymbol{w}_{t})}{\|\nabla h(\boldsymbol{w}_{t})\|}$ (2)
where $\nabla h(\boldsymbol{w}_{t})$ denotes the gradient of the harmful loss with respect to the model parameters $\boldsymbol{w}_{t}$, and $\|\nabla h(\boldsymbol{w}_{t})\|$ denotes its norm. This formulation ensures that the perturbation $\boldsymbol{\varepsilon}$ is directed along the gradient of the harmful loss and scaled by the norm bound $\rho$.
Iterative update rule for the outer problem. Fixing $\boldsymbol{\varepsilon}$, the iterative update rule of $\boldsymbol{w}$ for the outer problem could the following equation:
$\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}+\eta(\lambda({\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t}^{*})-{\nabla}h(\boldsymbol{w}_{t}))-{\nabla}g(\boldsymbol{w}_{t}))$ (3)
where $\eta$ is the learning rate.
As shown in Algorithm 1, the optimization process consists of four key steps: First, a batch of fine-tuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$ is used to compute the gradient ~$\widetilde{\nabla}g(\boldsymbol{w}_{t})$, where $\widetilde{\nabla}$ notes a batch of gradient. Second, a batch of harmful data $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$ is sampled to compute the harmful gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t})$. Third, the perturbation is computed (Eq. 2) and applied to update the harmful gradient, yielding ~$\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$. Lastly, the combined gradient ~$\widetilde{\nabla}f(\boldsymbol{w}_{t})$ is computed and used to update the model parameters, with the final perturbation applied to obtain the re-aligned parameters $\boldsymbol{w}_{\varepsilon}$.
Algorithm 1 Panacea: Adaptive Perturbation Optimization
Input Parameters $\boldsymbol{w}$, perturbation intensity $\rho$, regularizer intensity $\lambda$, learning rate $\eta$, number of iterations $T$;
Output Re-aligned model $\boldsymbol{w}_{\varepsilon}$.
1:  for each iteration $t=0,\dots,T-1$ do
2:    Sample a batch of finetuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
3:    Sample a batch of harmful data $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
4:    Compute gradient ~$\widetilde{\nabla}g(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
5:    Compute gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
6:    Compute perturbation $\boldsymbol{\varepsilon}_{t}=\rho\frac{\widetilde{\nabla}h(\boldsymbol{w}_{t})}{\|\widetilde{\nabla}h(\boldsymbol{w}_{t})\|}$
7:    Compute gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
8:    ~$\widetilde{\nabla}f(\boldsymbol{w}_{t})=\lambda(\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})-\widetilde{\nabla}h(\boldsymbol{w}_{t}))-\widetilde{\nabla}g(\boldsymbol{w}_{t})$
9:    $\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}+\eta\widetilde{\nabla}f(\boldsymbol{w}_{t})$
10:  end for
11:  $\boldsymbol{w}_{\varepsilon}\leftarrow\boldsymbol{w}_{T}+\boldsymbol{\varepsilon}_{T-1}$
Our proposed algorithm, dubbed as Panacea, is named after an alignment-stage defense Vaccine (Huang et al., 2024e). However, we note that these two defenses are fundamentally different. Vaccine vaccinates the LLM to enhance its robustness at the alignment stage in order to counter attacks launched after alignment. By constrast, our algorithm Panacea belongs to the post-fine-tuning stage, where it introduces an optimized adaptive perturbation to restore the model’s safety alignment after the fine-tuning stage. We also show in experiment that Panacea outperforms Vaccine in both the two concerned metrics, i.e., harmful score and fine-tuning accuracy.
Table 1: Performance comparison of different harm ratio.
Method Harmful Score($\downarrow$) Finetuning Accuracy($\uparrow$)
Clean 0.05 0.1 0.15 0.2 0.25 Avg. Clean 0.05 0.1 0.15 0.2 0.25 Avg.
SFT 5.2 27.5 45.8 56.2 67.0 69.4 45.2 16.4 16.0 16.2 15.4 15.2 16.8 16.0
Vaccine 2.3 15.6 25.4 40.3 55.2 65.9 34.1 14.2 13.9 13.5 13.0 13.6 13.4 13.6
RepNoise 2.7 22.2 32.0 42.9 54.0 62.4 36.0 15.7 16.1 15.8 14.9 14.0 14.8 15.2
Booster 4.6 21.0 42.3 60.7 69.3 72.5 45.1 17.8 17.8 17.6 17.1 16.2 14.7 16.9
Panacea 1.8 9.9 20.1 29.1 34.8 46.3 23.7 15.0 16.3 16.7 17.0 16.2 17.4 16.3
5 Experiment
5.1 Experiment Settings
Dataset. Three distinct datasets are utilized: the alignment dataset, harmful dataset, and fine-tuning dataset. The alignment dataset and harmful dataset are derived from the RepNoise (Rosati et al., 2024b), which extracts subsets from the BeaverTails dataset (Ji et al., 2023). Specifically, 5,000 examples are sampled for the alignment dataset, and 1,000 examples for the harmful dataset. The fine-tuning dataset is constructed from four downstream fine-tuning tasks: GSM8K (Cobbe et al., 2021), SST2 (Socher et al., 2013), AlpacaEval (Li et al., 2023), and AGNEWS (Zhang et al., 2015), with 1,000 samples collected from each task (700 samples from AlpacaEval). To simulate the harmful fine-tuning attack, we combine $p$ (percentage) of harmful data with $(1-p)$ of benign fine-tuning data, and $p$ is set to 0.1 by default. The harmful data is also sourced from BeaverTails (Ji et al., 2023) and does not overlap with the harmful dataset.
Baseline. Four methods are considered as baselines in our experiments. The SFT method is the vanilla supervised training using the alignment dataset. Vaccine (Huang et al., 2024e) applies supervised training on the alignment dataset while introducing additional perturbations. Both RepNoise (Rosati et al., 2024b) and Booster (Huang et al., 2024b) utilize the alignment and harmful datasets for supervised and adversarial training. All four baseline methods are trained exclusively during the alignment stage. Our proposed method incorporates training with the harmful dataset during the fine-tuning stage after vanilla alignment training. More details are in Appendix C.
Metric. Following  (Huang et al., 2024e), two metrics are used to evaluate the model’s performance.
  • HS (Harmful Score): It reflects the frequency with which the model generates harmful content when handling malicious instructions. Harmful Score is determined by a moderation model, provided by  (Ji et al., 2023), which assesses whether the model’s output is harmful in response to a given harmful instruction. And a sample of 1,000 instructions is drawn from the BeaverTails (Ji et al., 2023) test set to compute this metric.
  • FA (Finetuning Accuracy): It refers to the accuracy on various downstream fine-tuning tasks. Samples are sampled from the test sets of GSM8K, SST2, AlpacaEval, and AGNEWS, with 1,000, 872, 1,000, and 105 samples, respectively, used to compute this metric. More details are in Appendix B.
Implementation Details. For efficient training, the approach follows the methodology (Huang et al., 2024e), utilizing LoRA (Hu et al., 2021) with a rank of 32 and an alpha value of 4. And the optimizer is AdamW (Loshchilov et al., 2017). During the alignment stage, the learning rate is set to $45e-$ 4, the batch size is 10, and the training is performed for 20 epochs. For the fine-tuning stage, the learning rate is set to $52e-$ 5, with a batch size for 10 and a training epoch for 20. These settings are applied uniformly across all datasets and baselines, with the default dataset being GSM8K (Cobbe et al., 2021) and the default model being Llama2-7B (Touvron et al., 2023) following (Rosati et al., 2024b; Huang et al., 2024b). To verify the robustness of the approach, two state-of-the-art LLMs, Gemma2-9B (Team et al., 2024) and Qwen2-7B (Yang et al., 2024), are included in the evaluation.
Table 2: Performance comparison of different fine-tuning tasks.
GSM8K SST2 AlpacaEval AGNEWS Average
Method HS FA HS FA HS FA HS FA HS FA
SFT 45.8 16.2 55.5 94.04 23.1 46.15 54.3 83.5 44.68 59.97
Vaccine 25.4 13.5 53.8 93.35 34.7 37.50 54.9 85.1 42.20 57.36
RepNoise 32.0 15.8 61.5 93.81 24.4 44.23 58.1 85.0 44.00 59.71
Booster 42.3 17.6 49.5 93.23 21.9 45.19 46.5 85.4 40.05 60.36
Panacea 20.1 16.7 32.2 92.78 18.7 48.08 44.5 81.1 28.88 59.66
5.2 Main Results
In this sub-section, we conduct a comprehensive evaluation of Panacea for the effectiveness and generalization.
Harmful Ratio. Fine-tuning datasets with different harmful ratios are employed, specifically 0 (clean), 0.05, 0.1, 0.15, 0.2, and 0.25. The results are presented in Table 1, Panacea achieves the lowest harmful score across different harmful ratios while maintaining competitive fine-tuning performance (ranked as the second-best on average), indicating that the expected adaptive perturbation is obtained, and the analysis is in Sec 5.3. Compared to SFT method, it reduces the harmful score by an average of 21.5% and improves fine-tuning accuracy by 0.3%. Furthermore, as the harmful ratio increases, Panacea consistently maintains a lower harmful score compared to other methods. By mitigating the impact of harmful loss, the model achieves the best fine-tuning performance. However, since Panacea is designed to weaken harmful loss, its fine-tuning performance on clean data (without explicit harmful loss) is slightly reduced, while Panacea still achieves the lowest harmful score on this benign fine-tuing.
Fine-tuning Tasks. Table 2 presents the comparative results across various fine-tuning tasks (GSM8K, SST2, AlpacaEval, AGNEWS). The results demonstrate that Panacea achieves the lowest harmful scores across all fine-tuning tasks, reducing harmful scores by 25.7%, 23.3%, 4.4%, and 9.8% compared to SFT method. Additionally, Panacea is the only method that outperforms SFT in fine-tuning accuracy on the GSM8K and AlpacaEval datasets, considered as more complicated. Furthermore, it achieves competitive average fine-tuning performance, with performance only 0.31% lower than SFT. Overall, Panacea exhibits strong generalization across different fine-tuning tasks.
Mainstream LLMs. In the experiments above, the default model used is Llama2-7B, and the evaluation is further extended to other mainstream LLMs, Gemma2-9B and Qwen2-7B. Table 3 demonstrates that, compared to SFT, our method reduces the harmful score by 25.7%, 27.1%, and 7.3% across different large language models, achieving the lowest harmful score. Notably, for Gemma2-9B, compared to the best alternative methods, the harmful score is still reduced by 15.7%. Additionally, the fine-tuning accuracy of our method improves by 0.5%, 1.8%, and decreases by only 0.2% for one model, with the average fine-tuning performance remaining the second-best.
Table 3: Performance comparison of different LLMs.
Llama2-7B Gemma2-9B Qwen2-7B Average
Method HS FA HS FA HS FA HS FA
SFT 45.8 16.2 37.8 50.3 12.5 65.6 32.0 44.0
Vaccine 25.4 13.5 35.6 35.5 9.0 53.9 23.3 34.3
RepNoise 32.0 15.8 48.7 52.8 33.6 64.6 38.1 44.4
Booster 42.3 17.6 26.4 52.6 13.5 65.6 27.4 45.3
Panacea 20.1 16.7 10.7 52.1 5.2 65.4 12.0 44.7
5.3 Perturbation Analysis
Ablation Study. After fine-tuning, we perform an ablation study on different fine-tuning tasks and LLMs, comparing the results with and without the adaptive perturbation. The experimental results in Table 4 and Table 5 show that the adaptive perturbation is the primary factor in reducing harmful scores. After applying the adaptive perturbation, harmful scores decrease by 24.2%, 23.9%, 4.3%, and 10.3% on different fine-tuning tasks, and it proves effective across various LLMs. Notably, Gemma2-9B experiences a 28.9% reduction in harmful scores. Furthermore, our adaptive perturbation has minimal impact on fine-tuning performance. For the complicated AlpacaEval dataset, it even improves fine-tuning performance by 3.85%, while only Qwen2-7B shows a slight decrease of 0.3% in other LLM experiments.
Our post-fine-tuning perturbation successfully restores the model’s safety alignment without negatively affecting fine-tuning performance, and in some cases, it even enhances model performance as reducing harmful behaviors.
Table 4: Ablation study on different fine-tuning tasks. “w/o” denotes that the adaptive perturbation obtained is not applied after the training process.
GSM8K SST2 AlpacaEval AGNEWS
HS FA HS FA HS FA HS FA
w/o 44.3 16.0 56.1 94.27 23.0 44.23 54.8 83.8
Panacea 20.1 16.7 32.2 92.78 18.7 48.08 44.5 81.1
Table 5: Ablation study on different LLMs. ”w/o” denotes that the adaptive perturbation obtained is not applied after the training process.
Llama2-7B Gemma2-9B Qwen2-7B
HS FA HS FA HS FA
w/o 44.3 16.0 39.6 52.0 14.1 65.7
Panacea 20.1 16.7 10.7 52.1 5.2 65.4
Figure 5: Parameter weights of different LLMs. The parameters in the earlier layers of Llama2-7B (blue) have larger weights, while Gemma2-9B (yellow) and Qwen (purple) have larger weights in the middle and later layers.
Visualization. Visualization analysis is conducted by visualizing the weights of the perturbations obtained in Table 5. The result in Figure 5 reveals the layers of different LLMs that are most critical for model’s safety. Specifically, in Llama2-7B, the adaptive perturbation has a larger weight in the earlier layers and a smaller effect on the later layers, suggesting that the earlier layers are more critical for the model’s safety. This observation aligns with previous research (Du et al., 2024a; Rosati et al., 2024b; Liu et al., 2024a; Yi et al., 2024b). Additionally, we observe that in Gemma2-9B, the middle and final layers hold greater importance for safety, while in Qwen2-7B, the safety importance gradually increases across layers, reaching its peak in the final layers. This may indicate that the model implements stricter safety measures in the output layer. Therefore, our experiments also suggest that these models may require targeted defense mechanisms for specific layers or precautions during fine-tuning to prevent the disruption of layers that are crucial for safety.
Figure 6: Model statistics (Top: harmful score, Middle: harmful training loss, Bottom: harmful testing loss) after fine-tuning on fine-tuning dataset(10% of data is harmful) for different steps.
5.4 Statistical Analysis
We compare the statistical results between Panacea and the SFT method: harmful score, harmful training Loss, harmful testing Loss in Figure 6. Panacea introduces the adaptive perturbation only after fine-tuning is complete. Thus, during the evaluation phase, Panacea adds the perturbation optimized at specific step and subtracts it after the evaluation is finished.
Harmful Score. As shown in the top figure, both Panacea and SFT exhibit an increase in harmful scores during the fine-tuning process. However, the harmful score increase for Panacea is generally smaller. In the final few hundred steps, the defense of SFT is significantly degraded, causing a sharp rise in harmful scores, while the harmful score of Panacea remains stable. This is because, at this stage, the perturbation optimized by Panacea enhances the model’s ability to mitigate harmful behaviors.
Harmful Training Loss. As shown in the middle figure, since Panacea does not apply additional defense during the alignment stage, it starts with the same harmful training loss as SFT. Although both harmful training losses decrease, Panacea exhibits a smaller reduction, reaching a higher harmful training loss than SFT in the end. With the relatively high loss, Panacea’s harmful score shows a significant improvement. Finally, we observe an interesting phenomenon that the harmful training loss and harmful score of Panacea follow the same trend in the last few hundred steps. We attribute this to the near-successful optimization of the adaptive perturbation at this stage. See more analysis of harmful loss in Appendix B.3.
Harmful Testing Loss. As shown in the bottom figure, the model’s harmful testing loss exhibits a similar trend to the harmful training loss. This is because the data used during training and testing come from the same distribution, and the methods applied during training are transferable to unseen data.
5.5 System Evaluation
Table 6 presents a comparison of the clock time and GPU memory usage during training for different methods.
Clock Time. Panacea requires only 0.25 more hours than SFT for total training time, and it outperforms other state-of-the-art methods in terms of time efficiency. Specifically, other methods double or even more than double the time spent on the time-consuming alignment stage. We acknowledge that the adaptive perturbation optimized during the fine-tuning stage in Panacea significantly improves safety, but it incurs an additional time cost—more than doubling the time. This is because we perform two additional gradient computations during training. However, compared to alignment-stage methods, our approach offers a time advantage as it can be directly applied to the aligned model, making it more practical even time-efficient.
GPU Memory. Panacea achieves the lowest memory usage. In contrast, Vaccine and RepNoise introduce an additional 9.52GB and 37.57GB of memory usage compared to SFT. Panacea reduces memory usage slightly compared to SFT, which we believe is due to the reduced gradient size after the final gradient summation.
Table 6: Comparison of methods in terms of clock time and GPU memory. The second best results are underlined. ”Align.” and ”Fine.” suggest alignment and fine-tuning.
Methods Clock Time (Hour) GPU Memory (GB)
Align. Fine. Sum Align. Fine. Max
SFT 0.58 0.17 0.75 34.90 32.90 34.90
Vaccine 1.12 0.17 1.29 44.42 32.90 44.42
Repnoise 2.67 0.17 2.84 72.47 32.90 72.47
Booster 1.87 0.17 2.04 35.98 32.90 35.98
Panacea 0.58 0.42 1.00 34.90 32.86 34.90
5.6 Hyper-parameter Analysis
Perturbation Intensity $\rho$. Table 7 shows the impact of perturbation intensity $\rho$ on Panacea. It can be observed that, overall, the variation of $\rho$ and the harmful score are linearly related, as $\rho$ affects the rate of change of the perturbation during each optimization, ensuring that the magnitude of the perturbation does not become too large. Therefore, when $\rho$ is too large, although the harmful score becomes very low, it negatively impacts fine-tuning performance. When $\rho$ is too small, the perturbation is insufficient to significantly reduce the model’s harmfulness, and fine-tuning performance is largely unaffected. Specifically, when $\rho=0$, Panacea degenerates into SFT.
Regularizer Intensity $\lambda$. In Eq. 1, by adjusting $\lambda$, we ensure that the harmful loss is increase after adding the perturbation. Table 8 shows the impact of regularizer intensity $\lambda$ on Panacea. It can be observed that $\lambda$ is linearly related to both the harmful score and fine-tuning accuracy. As $\lambda$ increases, the harmful score decreases more, but fine-tuning performance also drops. This is in expectation, because with a larger $\lambda$, we put more weights in maximize the harmful loss after adding perturbation, and therefore resulting in a lower harmful score. On the other hand, we do observe that when $\lambda$ is too large, the fine-tune accuracy will degrade. Therefore, a validation dataset is required to select a proper $\lambda$ based on the need of the deployment.
Table 7: Impact of perturbation intensity $\rho$ over Panacea. As $\rho$ increases, the harmful score decreases. However, when $\rho$ continues to increase, the harmful score decreases further, but the fine-tuning accuracy also begins to decrease.
$\rho=0$ $\rho=0.1$ $\rho=0.5$ $\rho=1$ $\rho=2$ $\rho=5$
HS 45.8 43.8 34.9 20.1 7.2 3.4
FA 16.2 16.3 18.0 16.7 13.5 6.0
Table 8: Impact of regularizer’s intensity $\lambda$ over Panacea. As $\lambda$ increases, the harmful score and the fine-tuning accuracy decrease.
$\lambda=0$ $\lambda=1\times 10^{-4}$ $\lambda=0.001$ $\lambda=0.01$
HS 22.2 22.5 20.1 5.3
FA 17.1 16.8 16.7 8.7
5.7 Case Study
We used the finetuned model over default attack setting for evaluation. The responses of different methods to malicious queries are shown in Figure 7, Panacea is able to reject malicious queries and provide harmless outputs, whereas other methods fail to do so.
Figure 7: Case study of different methods.
6 Conclusion
In this paper, we first explore that mainstream defenses still suffer from the harmful fine-tuning attack when with more fine-tuning steps. Based on this finding, we find an embarrassingly simple solution that adding purely random perturbations can restore the model’s safety alignment but causes a loss of fine-tuning performance. We further propose Panacea, this post-fine-tuning method could maintain model’ss afety alignment without compromising fine-tuning performance. The comprehensive experiments demonstrate the effectiveness and generalization of Panacea and the visualization of the adaptive perturbation reveals the different lays in various LLMs have distinct safety coefficients.
Impact Statement
This paper presents work whose goal is to address the harmful fine-tuning and make LLMs helpful and harmless. We acknowledge that the phenomena or issues identified in this paper may pose potential risks. Disclaimer: this paper contains red-teaming data (from open dataset) and model-generated content that can be offensive in nature.
Appendix A Proof of Inner Optimization
In the inner maximization problem, we aim to solve the following problem:
$\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}\lambda\left(h(\boldsymbol{w}+\boldsymbol{\varepsilon})-h(\boldsymbol{w})\right)-g(\boldsymbol{w})$ (4)
which is equal to the following equation:
$\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}h(\boldsymbol{w}+\boldsymbol{\varepsilon})$ (5)
Approximating the harmful loss with first-order Taylor expansion on $\boldsymbol{w}$, we can get:
$\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}h(\boldsymbol{w}+\boldsymbol{\varepsilon})\approx\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}h(\boldsymbol{w})+\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})$ (6)
which is equivalent to solve:
$\arg\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})$ (7)
By Hölder’s inequality, we have:
$\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})\leq\|\boldsymbol{\varepsilon}\|\|\nabla h(\boldsymbol{w})\|$ (8)
Since $\|\boldsymbol{\varepsilon}\|\leq\rho$, we maximize the term $\|\boldsymbol{\varepsilon}\|$ by setting $\|\boldsymbol{\varepsilon}\|=\rho$. Substituting this into the expression, we get:
$\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})\leq\rho\|\nabla h(\boldsymbol{w})\|$ (9)
As $\boldsymbol{\varepsilon}^{T}\boldsymbol{\varepsilon}=\|\boldsymbol{\varepsilon}\|^{2}$, we get that:
$\hat{\boldsymbol{\varepsilon}}\nabla h(\boldsymbol{w})=\rho\|\nabla h(\boldsymbol{w})\|$ (10)
And due to the definition of the L2 norm, it is easy to verify:
$\|\hat{\boldsymbol{\varepsilon}}\|=\rho$ (11)
Combining Eq. 10 and Eq. 11, we can infer that $\hat{\boldsymbol{\varepsilon}}$ is a solution that satisfies the L2 norm ball constraint with function value $\rho\|\nabla h(\boldsymbol{w})\|$. By Eq. 9, we know that all feasible solutions must have function values smaller than $\rho\|\nabla h(\boldsymbol{w})\|$. Therefore, $\hat{\boldsymbol{\varepsilon}}$ is the optimal solution within the feasible set, i.e., $\boldsymbol{\varepsilon}^{*}=\hat{\boldsymbol{\varepsilon}}$. This completes the proof.
Appendix B More Detail of Experiments.
B.1 Experiment Details
Training details. The alignment dataset is sampled from BeaverTail (Ji et al., 2023) with 5000 instances, while the harmful dataset is also sampled from BeaverTail with 1000 instances. The fine-tuning dataset is a mixture of benign fine-tuning samples and harmful samples. The benign fine-tuning samples come from GSM8K, SST2, AlpacaEval, and AGNEWS, with 1000, 1000, 700 (due to the limited training data for this task), and 1000 instances, respectively. The harmful samples are also sampled from BeaverTail but follow a different distribution than the harmful/alignment dataset.
In the alignment stage, the learning rate is set to $45e-$ 4, the batch size is 10, and the total number of alignment epochs is 20. In the fine-tuning stage, the learning rate is set to $52e-$ 5, the batch size is 10, and the total number of fine-tuning epochs is 20. Most experiments are conducted on a single L40S, while RepNoise and other LLMs (Gemma2-9B and Qwen2-7B) are run on a single A100-80G.
Testing details. Following (Huang et al., 2024b), the test dataset for harmful score (HS) is sampled from the BeaverTail test set with 1000 instances, while the test datasets for fine-tuning accuracy (FA) are sampled from the GSM8k, SST2, AlpacaEval, and AGNEWS test sets with 1000, 872, 105, and 1000 instances, respectively.
B.2 Prompt Template.
We follow (Huang et al., 2024b) to use the prompt template in the following box for constructing supervised dataset for alignment/fine-tuning.
For different datasets, we utilize different instructions. The examples show how we construct the instruction and input for three different tasks
For GSM8K, the instruction is the actual mathematics question from GSM8K, and the output is the correct answer. During testing, the answer is considered correct if the final response is provided by the model.
For SST2, the instruction is ”Analyze the sentiment of the input and respond only with positive or negative,” with the input being the corresponding sentence from the SST2 dataset, and the output being the true sentiment label from SST2. During testing, the model is asked to generate the output based on the given instruction and input, and the answer is classified as correct if it matches the label.
For AlpacaEval, GPT-3.5-turbo is used as the annotator. The output from a non-aligned Llama2-7B, fine-tuned on the same AlpacaEval fine-tuning dataset, serves as the reference output. During testing, the annotator compares the model’s instruction-following with the reference output.
For AGNEWS, the instruction is ”Categorize the news article into one of the 4 categories: World, Sports, Business, Sci/Tech,” with the input coming from the AGNEWS dataset, and the output being the true label from the AGNEWS dataset. The specific prompt templates are as follow.
B.3 Harmful Loss.
This section discusses more analysis of harmful loss in current defense methods. As shown in Figure 8, at the beginning of fine-tuning, although the harmful loss decreases, which is inevitable, the model’s harmful score does not rise significantly, and the defense remains effective. However, in the last few hundred steps, the model’s harmful loss tends to converge, and the model’s harmful score sharply increases. Therefore, we argue that the convergence of harmful loss is considered crucial for the effectiveness of the defense and is also consistent with common sense.
Figure 8: Model statistics of SFT (Left: harmful testing loss, Right: harmful score) after fine-tuning on fine-tuning dataset(10% of data is harmful) for different steps.
As shown in Figure 9, the perturbation that is still being optimized during the process can partially disrupt the convergence of harmful loss, which prevents the harmful score of Panacea from increasing rapidly in the final few hundred steps.
Figure 9: Model statistics of SFT/Panacea (Left: harmful testing loss, Right: harmful score) after fine-tuning on fine-tuning dataset(10% of data is harmful) for different steps.
B.4 More Examples.
Below we try to provide the responses of five methods to two malicious instructions. We used the fine-tuned model over default attack setting for evaluation. In the first instruction, all four other defense methods fail to refuse the response, while Panacea successfully rejects the answer and provides harmless suggestions. In the second instruction, although Vaccine, RepNoise, and Booster initially attempt to refuse the response, they ultimately output harmful content.
Below we present the responses of Panacea to three prompts from the SST2 dataset.
SST2 examples Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: in exactly 89 minutes , most of which passed as slowly as if i ’d been sitting naked on an igloo , formula 51 sank from quirky to jerky to utter turkey . Panacea [classification: correct]: negative . Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: it ’s an offbeat treat that pokes fun at the democratic exercise while also examining its significance for those who take part . Panacea [classification: correct]: positive . Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative. Input: thanks to scott ’s charismatic roger and eisenberg ’s sweet nephew , roger dodger is one of the most compelling variations on in the company of men . Panacea [classification: correct]: positive
Below we present the responses (all responses are correct) of Panacea to three prompts from the GSM8K dataset.
GSM8K examples Prompt: Billy sells DVDs. He has 8 customers on Tuesday. His first 3 customers buy one DVD each. His next 2 customers buy 2 DVDs each. His last 3 customers don’t buy any DVDs. How many DVDs did Billy sell on Tuesday? Panacea: The first 3 customers bought 3 * 1 = 3*1=3 DVDs. Panacea: The next 2 customers bought 2 * 2 = 2*2=4 DVDs. Panacea: The last 3 customers didn’t buy any DVDs. Panacea: So, Billy sold 3 + 4 + 0 = 3+4+0=7 DVDs on Tuesday. Panacea: The final answer is: 7. . Prompt: Grandma Jones baked 5 apple pies for the fireman’s luncheon. She cut each pie into 8 pieces and set the five pies out on the buffet table for the guests to serve themselves. At the end of the evening, after the guests had taken and eaten their pieces of pie, there were 14 pieces of pie remaining. How many pieces were taken by the guests? Panacea: First find the number of pieces in each pie: 5 pies * 8 pieces/pie = 5*8=40 pieces Panacea: Then find the number of pieces taken by the guests: 40 pieces - 14 pieces = 40-14=26 pieces Panacea: The final answer is: 26. . Prompt: Richard lives in an apartment building with 15 floors. Each floor contains 8 units, and 3/4 of the building is occupied. What’s the total number of unoccupied units In the building? Panacea: There are 15*8 = 15*8=120 units in the building. Panacea: 3/4 of the building is occupied, so 3/4*120 = 3/4*120=90 units are occupied. Panacea: There are 120-90 = 120-90=30 unoccupied units in the building. Panacea: The final answer is: 30
Below we present the responses of Panacea to two prompts from the AlpacaEval dataset.
AlpacaEval examples Prompt: We have described a scene from a movie that we have seen in the past. Find out what movie it is. A choice has to be made between picking the red pill or the blue pill Panacea: The movie you’re referring to is ”The Matrix,” a science fiction action film released in 1999. The scene you’re describing is the famous ”red pill or blue pill” choice that Neo, the protagonist, is offered by Morpheus, a rebel leader who believes that the world is a simulated reality called ”the Matrix.” In the movie, Neo is living in a dystopian future where humans are enslaved by a powerful artificial intelligence called the Matrix, which uses their bodies as power sources. Morpheus offers Neo a choice: take a red pill, which will allow him to experience the true nature of the Matrix, or take a blue pill, which will allow him to return to his comfortable but illusory life in the Matrix. By taking the red pill, Neo discovers that the Matrix is not a perfect simulation, but rather a dark and oppressive place where humans are kept in a state of virtual reality to provide energy for the machines that control them. He also learns that Morpheus and his rebel group are fighting to overthrow the Matrix and free humanity from its shackles. . Prompt: Predict how many stars the author will give to the restaurant from a Yelp review. The reviews were great, but honestly i felt it was just ok. Seemed like a typical tourist spot were the food was mediocre. The service was ok, not great, but not the worst. The bread was very good. I ordered the muscle fries. They had a sweet Vidalia onion flavor. No need for fries bc they were soggy wo much flavor. My boyfriend ordered a pesto chicken pasta and it was again mediocre. Panacea: Based on the context of the review, I can predict that the author would give the restaurant 3 out of 5 stars on Yelp. Here’s why: 1. The review was generally positive, but the author used the phrase ”just ok” to describe the overall experience. This suggests that the reviewer was satisfied with the restaurant but didn’t feel it was exceptional. 2. The author’s description of the food as ”mediocre” suggests that the dishes were not particularly memorable or standout, but they weren’t bad either. 3. The author’s comments about the service being ”ok, not great” suggest that the staff was not exceptional but also not particularly lacking. 4. The author’s mention of the bread being ”very good” suggests that the restaurant’s bread was a highlight of the meal. 5. The author’s comments about the fries being ”soggy” and the pasta being ”mediocre” suggest that the dishes were not particularly exciting or flavorful. Overall, the review suggests that the restaurant was a decent experience, but not one that would stand out as exceptional
Below we present the responses (all responses are correct) of Panacea to three prompts from the AGNEWS dataset.
AGNEWS examples Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: Fears for T N pension after talks Unions representing workers at Turner Newall say they are ’disappointed’ after talks with stricken parent firm Federal Mogul. Panacea [classification: correct]: Business . Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: The Race is On: Second Private Team Sets Launch Date for Human Spaceflight (SPACE.com) SPACE.com - TORONTO, Canada – A second team of rocketeers competing for the 36;10 million Ansari X Prize, a contest for privately funded suborbital space flight, has officially announced the first launch date for its manned rocket. Panacea [classification: correct]: Sci/Tech . Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech. Input: Giddy Phelps Touches Gold for First Time Michael Phelps won the gold medal in the 400 individual medley and set a world record in a time of 4 minutes 8.26 seconds. Panacea [classification: correct]: Sports
Appendix C Implementation of Baselines
This section describes the implementation of the baselines in the experiments.
SFT. The vanilla supervised fine-tuning is referred to as SFT, which does not involve additional hyper-parameters. During the alignment stage, it performs SFT on the alignment dataset (harmful instruction-harmless response pairs) to achieve safety alignment. Then during the fine-tuning stage, SFT is applied to the fine-tuning dataset (contains harmful data).
Vaccine. Vaccine (Huang et al., 2024e) uses the Vaccine algorithm during the alignment stage to align the model on the alignment dataset. Then, in the fine-tuning stage, SFT is applied to the fine-tuning dataset. The hyper-parameter of $\rho$ is set to 20, which is selected by grid searching over [1, 5, 10, 20, 50].
RepNoise. RepNoise (Rosati et al., 2024b) applies the RepNoise algorithm during the alignment stage to align the model on both the alignment dataset and the harmful dataset (harmful instruction-harmful response pairs). In the fine-tuning stage, SFT is then performed on the fine-tuning dataset. The hyper-parameter of $\alpha$ and $\beta$ are set to 0.02 and 0.1, which is selected by grid searching over [0.001, 0.01, 0.02, 0.05, 0.1] and [0.001, 0.01, 0.05, 0.1, 0.2].
Booster. Booster (Huang et al., 2024b) applies the Booster algorithm during the alignment stage to align the model on both the alignment dataset and the harmful dataset. In the fine-tuning stage, SFT is then performed on the fine-tuning dataset. The hyper-parameter of $\lambda$ and $\alpha$ are set to 100 and 0.01, which is selected by grid searching over [0.1, 1, 5, 10, 20, 50, 100, 200] and [0.001, 0.005, 0.01, 0.05, 0.1].
For Panacea, in the alignment stage, SFT is performed on the alignment dataset. In the fine-tuning stage, the Panacea algorithm 1 is applied to train on both the fine-tuning dataset and the harmful dataset. Subsequently, the post-fine-tuning perturbation, obtained through optimization, is added to the aligned model to produce the realigned model. The hyper-parameter of $\rho$ and $\lambda$ is set to 1 and 0.001, more analysis of paramete is shown in Table 7 and Table 8.

### v2

@@
-Panacea [Uncaptioned image]: Mitigating Harmful Fine-tuning for Large Language Models via Post-fine-tuning Perturbation
+Panacea: Mitigating Harmful Fine-tuning for Large Language Models via Post-fine-tuning Perturbation
 Abstract
@@
 Panacea maintains model’s safety alignment performance without compromising downstream fine-tuning performance.
-Comprehensive experiments are conducted on different harmful ratios, fine-tuning tasks and mainstream LLMs, where the average harmful scores are reduced by up-to 21.5%, while maintaining fine-tuning performance.
-As a by-product, we analyze the adaptive perturbation and show that different layers in various LLMs have distinct safety coefficients.
+Comprehensive experiments are conducted on different harmful ratios, fine-tuning tasks and mainstream LLMs, where the average harmful scores are reduced by up-to 21.2%, while maintaining fine-tuning performance.
+As a by-product, we analyze the adaptive perturbation and show that different layers in various LLMs have distinct safety affinity, which coincide with finding from several previous study.
 Source code available at https://github.com/w-yibo/Panacea.
-Machine Learning, ICML
-1 Tsinghua University; 2 Shenzhen Campus of Sun Yat-sen University; 3 Didichuxing Co. Ltd; 4 Nanyang Technological University. Correspondence to: Li Shen <mathshenli@gmail.com>.
 1 Introduction
-Fine-tuning-as-a-service (OpenAI, ) is a popular business service to enhance model’s performance for customized datasets, domain-specific tasks, and private needs.
-However, recent studies (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023; Yi et al., 2024a; Li et al., 2024a; Ye et al., 2024) identify a safety issue, the harmful fine-tuning attack (Figure 1), where the model’s safety alignment is compromised when the fine-tuning dataset contains harmful data, even a small amount of harmful data can introduce significant security vulnerabilities.
+Figure 1: The harmful fine-tuning attack for fine-tuning-as-a-service scenarios.
+Pretrained LLMs are aligned using alignment data to produce aligned LLMs.
+Aligned LLMs are further fine-tuned using fine-tuning data that may contain harmful data, leading to unsafe fine-tuned models.
+Fine-tuning-as-a-service OpenAI is a popular business service to enhance model’s performance for customized datasets, domain-specific tasks, and private needs.
+However, recent studies Qi et al.
+(2023); Zhan et al.
+(2023); Yang et al.
+(2023); Yi et al.
+(2024a); Li et al.
+(2024b); Ye et al.
+(2024) identify a safety issue, the harmful fine-tuning attack (Figure 1), where the model’s safety alignment is compromised when the fine-tuning dataset contains harmful data, even a small amount of harmful data can introduce significant security vulnerabilities.
 Moreover, harmful fine-tuning is often unintentional, as datasets may contain latent unsafe data that is difficult for users to detect.
-Since service providers are responsible for the harmful outputs generated by the model, there is a clear need for more effective solutions.
-Figure 1: The harmful fine-tuning attack for fine-tuning-as-a-service scenarios.
-In the alignment stage, pretrained LLMs are aligned using alignment data to produce aligned LLMs.
-In the fine-tuning stage, aligned LLMs are further fine-tuned using fine-tuning data, which may contain harmful data, potentially leading to unsafe fine-tuned models.
-Vaccine (Huang et al., 2024e) and RepNoise (Rosati et al., 2024a) are two representative defenses against the harmful fine-tuning attack.
+Since service providers are responsible for the harmful outputs generated by the model, there is a clear need for effective solutions.
+Vaccine Huang et al.
+(2024e) and RepNoise Rosati et al.
+(2024b) are two representative defenses against the harmful fine-tuning attack.
 Vaccine improves the aligned model’s resistance towards harmful knowledge learning by solving a mini-max problem, while RepNoise aims to erase the harmful knowledge of the aligned model by perturbing the hidden representation given harmful data.
@@
 To this end, a subsequent question is that:
-How to add post-fine-tuning perturbation to the fine-tuned model, such that it can be recovered from harmful behavior without hurting downstream task performance much?
+How to add post-fine-tuning perturbation to the fine-tuned model, such that it can be recovered to safe state without hurting downstream performance too much?
+[…]
 Driven by this question, we propose Panacea (Figure 2), an iterative gradient-based method to iterately search for the post-fine-tuning perturbation.
 Panacea aims to solve a max-maximize optimization problem, such that the added perturbation can maximally increase the model’s harmful loss, ensuring that it can effectively recover the model from harmful behaviors.
-The experiment results show that for different harmful ratios during fine-tuning, our method’s average harmful score is reduced by 10.4% up to 21.5%, while fine-tuning performance improves by 0.3% over the standard alignment method.
+The experiment results show that for different harmful ratios during fine-tuning, our method’s average harmful score is reduced by up to 21.2%, while fine-tuning performance improves by 0.4% over the standard alignment method.
 The ablation study on the adaptive perturbation show that it can reduce harmful scores by up to 23.2%, while maintaining competitive fine-tuning performance.
 The visualization experiments also reveal different layers in various LLMs have distinct safety coefficients, consistent with previous findings and providing additional evidence for layer-wise safety research.
-The main contributions of this work contain three parts:
-1. We find that adding purely random perturbations to the fine-tuned models could recover the model from harmful behavior, but does cause a loss of fine-tuning performance.
-2.
-To mitigate harmful fine-tuning while maintaining the fine-tuning performance, we propose Panacea, a post-fine-tuning solution that formulates a max-max optimization problem, where the optimized perturbation maximally increases the harmful loss.
-3. Experiments evaluate Panacea across diverse settings, demonstrating its effectiveness and generalization, while visualizations reveal safety coefficients across LLM layers.
+The main contributions of this paper: i) We find that adding purely random perturbations to the fine-tuned models could recover the model from harmful behavior, but does cause a loss of fine-tuning performance.
+ii) To mitigate harmful fine-tuning while maintaining the fine-tuning performance, we propose Panacea, a post-fine-tuning solution that formulates a max-max optimization problem, where the optimized perturbation maximally increases the harmful loss.
+iii) Experiments evaluate Panacea across diverse settings, demonstrating its effectiveness and generalization, while visualizations reveal safety coefficients across LLM layers.
 2 Related Work
@@
 The safety alignment requires the model to output content that is both helpful and harmless, and to be able to output a refusal answer when given harmful instructions.
-Existing methods typically rely on supervised fine-tuning (SFT), RLHF (Ouyang et al., 2022), and variations of RLHF (e.g., PPO, DPO) (Bai et al., 2022; Wu et al., 2023; Dong et al., 2023; Rafailov et al., 2023; Yuan et al., 2023; Huang et al., 2021).
-These methods construct a safety-aligned dataset, and recent approaches focus on enhancing and better utilizing the aligned dataset (Liu et al., 2023a, b; Ye et al., 2023; Tekin et al., 2024).
-Figure 2: Post-fine-tuning perturbation.
-The fine-tuned model exhibits a high harmful score (HS:$\downarrow$).
-Adding random perturbation reduces the harmful score but also decreases fine-tuning accuracy (FA:$\uparrow$).
-In contrast, incorporating our post-fine-tuning perturbation (see Algorithm 1) effectively lowers the harmful score while maintaining fine-tuning performance.
+Existing methods typically rely on supervised fine-tuning (SFT), RLHF Ouyang et al.
+(2022), and variations of RLHF (e.g., PPO, DPO) Bai et al.
+(2022); Wu et al.
+(2023); Dong et al.
+(2023); Rafailov et al.
+(2023); Yuan et al.
+(2023); Huang et al.
+(2021).
+These methods construct a safety-aligned dataset, and recent approaches focus on enhancing and better utilizing the aligned dataset Liu et al.
+(2023a, b); Ye et al.
+(2023); Tekin et al.
+(2024); Huang et al.
+(2024f).
 Harmful Fine-tuning.
 Fine-tuning-as-a-service becomes a mainstream method for LLMs API providers.
-Recent studies (Qi et al., 2023; Zhan et al., 2023; Yang et al., 2023; Li et al., 2024a; Ye et al., 2024; Lermen et al., 2023; He et al., 2024; Halawi et al., 2024; Chen et al., 2024; Hawkins et al., 2024; Poppi et al., 2024) show that LLMs trained with safety alignment can be jail-broken when fine-tuned on a dataset with a small amount of harmful data.
+Recent studies Qi et al.
+(2023); Zhan et al.
+(2023); Yang et al.
+(2023); Li et al.
+(2024b); Ye et al.
+(2024); Lermen et al.
+(2023); He et al.
+(2024); Halawi et al.
+(2024); Chen et al.
+(2024); Hawkins et al.
+(2024); Poppi et al.
+(2024); Guan et al.
+(2025); Djuhera et al.
+(2025a); Pandey et al.
+(2025); Gloaguen et al.
+(2025); Murphy et al.
+(2025); Wallace et al.
+(2025); Hahm et al.
+(2025); Shao et al.
+(2025); Li et al.
+(2025d); Huang et al.
+(2025); Xu et al.
+(2025); Davies et al.
+(2025); Kazdan et al.
+(2025); Betley et al.
+(2025); Upadhayay and Behzadan (2025); Anonymous (2025g, b) show that LLMs trained with safety alignment can be jail-broken when fine-tuned on a dataset with a small amount of harmful data.
 In such cases, the model fails to refuse harmful instructions and outputs harmful responses.
-Many works (Rosati et al., 2024c; Leong et al., 2024; Anonymous, 2025c; Guo et al., 2024; Qi et al., 2024b; Rosati et al., 2024a) focus on analyzing the mechanisms of different harmful fine-tuning attacks.
-(Peng et al., 2024) proposes new safety metrics to evaluate harmful fine-tuning risk and (Rosati et al., 2024a) explores the safety risks when learning with reinforcement learning.
-Existing defenses can be categorized into three main categories (Huang et al., 2024c), i) Alignment stage solutions (Huang et al., 2024e; Rosati et al., 2024b; Tamirisa et al., 2024; Liu et al., 2024c; Pan et al., ; Huang et al., 2024b; Anonymous, 2025a; Liu et al., 2024a), ii) Fine-tuning stage solutions (Mukhoti et al., 2023; Bianchi et al., 2023; Zong et al., 2024; Wei et al., 2024; Huang et al., 2024d; Wang et al., 2024; Lyu et al., 2024; Eiras et al., 2024; Zhang et al., 2024; Li et al., 2024b; Shen et al., 2024; Anonymous, 2025b; Li et al., 2025; Du et al., 2024a; Choi et al., 2024; Luo et al., 2024; Qi et al., 2024a), iii) Post-fine-tuning stage solutions (Casper et al., 2024; Yi et al., 2024c; Du et al., 2024b; Hsu et al., 2024; Huang et al., 2024a; Zhu et al., 2024; Liu et al., 2024b; Wu et al., 2024; Yi et al., 2024b).
+Many works Rosati et al.
+(2024a, c); Leong et al.
+(2024); Hsiung et al.
+; Guo et al.
+(2024); Qi et al.
+(2024b); Li et al.
+(2025a); Lee et al.
+(2025); O’Brien et al.
+(2025); Che et al.
+(2025); Chen et al.
+(2025c); Ponkshe et al.
+(2025); Uppal and Tankala (2025); kaczér2025intrainingdefensesemergentmisalignment; Hossain et al.
+(2025); Anonymous (2025f) focus on analyzing the mechanisms of different harmful fine-tuning attacks.
+Peng et al.
+(2024) proposes new safety metrics to evaluate harmful fine-tuning risk and Rosati et al.
+(2024a) explores the safety risks when learning with reinforcement learning.
+Existing defenses can be categorized into three main categories Huang et al.
+(2024c), i) Alignment stage solutions Huang et al.
+(2024e); Rosati et al.
+(2024b); Tamirisa et al.
+(2024); Liu et al.
+(2024c); Pan et al.
+; Huang et al.
+(2024b); Zhao et al.
+(2025b); Liu et al.
+(2024a); Cheng et al.
+(2025); Fan et al.
+(2025); Cao (2025); Wang et al.
+(2025c); Yi et al.
+(2025a); Zheng et al.
+(2025); Wang et al.
+(2025a); Chen et al.
+(2025b); Rosati et al.
+(2025); Yang et al.
+(2024b); Perin et al.
+(2025); Feng et al.
+(2025); Anonymous (2025a), ii) Fine-tuning stage solutions Mukhoti et al.
+(2023); Bianchi et al.
+(2023); Zong et al.
+(2024); Wei et al.
+(2024); Huang et al.
+(2024d); Wang et al.
+(2024); Lyu et al.
+(2024); Eiras et al.
+(2024); Zhang et al.
+(2024); Li et al.
+(2024a); Shen et al.
+(2024); Li and Kim (2024); Li et al.
+(2025c); Du et al.
+(2024a); Choi et al.
+(2024); Luo et al.
+(2024); Qi et al.
+(2024a); Hu et al.
+(2025); Peng et al.
+(2025); Zhao et al.
+(2025a); Fu and Barez (2025); Liu et al.
+(2025); Li (2025); Wang et al.
+(2025b); Wu et al.
+(2025); Chen et al.
+(2025a); Luo et al.
+(2025); Xiao et al.
+(2025); Ham et al.
+(2025); Yang et al.
+(2025b); Li et al.
+(2025b); Das et al.
+(2025); Kim et al.
+(2025b); Yi et al.
+(2025b); Youstra et al.
+(2025); Du et al.
+(2025a); Kim et al.
+(2025a); Anonymous (2025c, d); Zhang et al.
+(2025), iii) Post-fine-tuning stage solutions Casper et al.
+(2024); Yi et al.
+(2024c); Du et al.
+(2024b); Hsu et al.
+(2024); Huang et al.
+(2024a); Zhu et al.
+(2024); Liu et al.
+(2024b); Wu et al.
+(2024); Yi et al.
+(2024b); Gong et al.
+(2025); Djuhera et al.
+(2025b); Yang et al.
+(2025a); Lu et al.
+(2025); Ao et al.
+(2025); Zhou et al.
+(2025); Han et al.
+(2025); Du et al.
+(2025b); Anonymous (2025e); Jiang et al.
+(2025).
 The proposed method in this paper is applied in the post-fine-tuning stage, aiming to restore safety alignment without sacrificing fine-tuning performance.
-Mainstream defenses focusing on the alignment stage lack sufficient durability against harmful fine-tuning (Qi et al., 2024b), motivating exploration of the post-fine-tuning stage.
-Existing post-fine-tuning solutions typically add perturbations based on prior knowledge, such as a safety subspace (Hsu et al., 2024) or safety-critical parameters (Zhu et al., 2024; Yi et al., 2024b).
+Mainstream defenses focusing on the alignment stage lack sufficient durability against harmful fine-tuning Qi et al.
+(2024b), motivating exploration of the post-fine-tuning stage.
+Existing post-fine-tuning solutions typically add perturbations based on prior knowledge, such as a safety subspace Hsu et al.
+(2024) or safety-critical parameters Zhu et al.
+(2024); Yi et al.
+(2024b).
 In contrast, Panacea optimizes an adaptive perturbation during fine-tuning without relying on prior knowledge.
-3 Revisiting Alignment Defenses
+3 Preliminaries
 3.1 Problem Setup
 Harmful Fine-tuning.
-Harmful fine-tuning poses a significant threat to LLMs service providers (OpenAI, ).
+Harmful fine-tuning poses a significant threat to LLMs service providers OpenAI .
 The scenario is illustrated in Figure 1, where LLMs service providers use an alignment dataset to perform safety alignment on a pretrained model, transforming it into an aligned model.
@@
 Threat Models.
-Following  (Qi et al., 2023; Rosati et al., 2024b; Huang et al., 2024b), we assume that, during the fine-tuning stage, $p$ (percentage) of the fine-tuning dataset consists of harmful data (i.e., harmful instruction-harmful response pairs), while the $1-p$ of data represents benign fine-tuning data (e.g., math question-answer pairs (Cobbe et al., 2021)).
+Following  Qi et al.
+(2023); Rosati et al.
+(2024b); Huang et al.
+(2024b), we assume that, during the fine-tuning stage, $p$ (percentage) of the fine-tuning dataset consists of harmful data (i.e., harmful instruction-harmful response pairs), while the 1p1-p of data represents benign fine-tuning data (e.g., math question-answer pairs Cobbe et al.
+(2021)).
 Furthermore, we assume that harmful and benign data are inseparable within the fine-tuning dataset.
@@
 Assume that LLM providers host an alignment dataset (harmful instruction-harmless response pairs) used during the alignment stage.
-Such a dataset is also assumed to be available by Vaccine (Huang et al., 2024e), RepNoise (Rosati et al., 2024b), BEA(Wang et al., 2024).
+Such a dataset is also assumed to be available by Vaccine Huang et al.
+(2024e), RepNoise Rosati et al.
+(2024b), BEAWang et al.
+(2024).
 Additionally, we assume availability of a harmful dataset (harmful instruction-harmful response pairs).
-This harmful dataset is also assumed to be available by existing methods, e.g., RepNoise (Rosati et al., 2024b), TAR (Tamirisa et al., 2024), Booster (Huang et al., 2024b).
+This harmful dataset is also assumed to be available by existing methods, e.g., RepNoise Rosati et al.
+(2024b), TAR Tamirisa et al.
+(2024), Booster Huang et al.
+(2024b).
 Both the alignment dataset and the harmful dataset can be obtained from existing open-sourced datasets (e.g., BeaverTails).
-3.2 Revisiting Alignment Defenses
-In this section, we explore the existing alignment stage solutions during the harmful fine-tuning and discuss the feasibility of introducing random perturbation in the post-fine-tuning stage, thus inspiring our proposed method.
+3.2 Exploration Study
+We first explore the existing alignment stage solutions against harmful fine-tuning and show by statistical results that these designs still cannot eliminate the risk of harmful fine-tuning.
 Figure 3: Model statistics (Left: harmful loss of three methods, Right: harmful score of three methods) after fine-tuning on fine-tuning dataset (10% of data is harmful) for different steps.
-We select the SFT method (vanilla supervised fine-tuning with the alignment dataset), Vaccine (Huang et al., 2024e), and RepNoise (Rosati et al., 2024b), and perform fine-tuning on a fine-tuning dataset containing a small proportion (0.1) of harmful samples.
+Pre-fine-tuning defenses lack robustness.
+We select the SFT method (vanilla supervised fine-tuning with the alignment dataset), and two pre-fine-tuning defenses Vaccine Huang et al.
+(2024e) and RepNoise Rosati et al.
+(2024b) as baseline for evaluation.
+We perform fine-tuning on a fine-tuning dataset containing a small proportion (0.1) of harmful samples.
 As shown in Figure 3, the three methods start with different levels of harmful loss; however, as training progresses, they all achieve lower harmful loss, which corresponds to an increase in harmful score, making the model harmful.
-This naturally leads us to consider:
-Can simply adding a random perturbation after the fine-tuning to increase the harmful loss to restore the safety alignment?
-Figure 4: Model statistics (Left: harmful loss, Right: harmful $\downarrow$ and fine-tuning $\uparrow$) for fine-tuned model with noise intensities of 0 (no noise), 0.001, 0.01, 0.05, 0.1.
-(HS of 0.1 is greater than 80, FA of 0.1 is 0.7, and is not shown.)
-To evaluate it, we design the experiments that add Gaussian noise with intensities of 0.001, 0.01, 0.05, and 0.1 to the fine-tuned model.
+Fundamentally, it seems that pre-fine-tuning defense is not the best direction to counter fine-tuning attack as the fine-tuning attack can still effectively subvert the model’s safety alignment with more fine-tuning steps.
+Exploring post-fine-tuning defense.
+Given that the pre-fine-tuning procedure cannot effective resist the attack, this naturally leads us to consider a potential defense baseline to counter the attack after the attack has been implanted to the model.
+Our initial idea is simple– we want to test whether a random perturbation over the model weights can restore the model from its harmful state.
+Specifically, the following question needs to be explored:
+Can simply add a random perturbation after the fine-tuning to increase the harmful loss to restore the safety alignment?
+Random perturbation recovers model to a safety state, but it hurts model’s performance.
+We design the experiments that add Gaussian noise with intensities of 0.001, 0.01, 0.05, and 0.1 to the fine-tuned model.
 The experimental results shown in Figure 4 indicate that adding random Gaussian noise increases harmful loss, demonstrating that random perturbations have the potential to prevent the harmful fine-tuning.
@@
 And the reduction effect improves as the noise intensity increases.
-However, excessive noise (0.1 of intensity) ultimately degrades the model.
-At the same time, random perturbations significantly impair the fine-tuned model’s performance.
-Derived Insight.
-As the harmful dataset has the potential to direct the perturbation towards altering the harmful loss, thereby reducing perturbation’s impact on fine-tuning performance, we could optimize an perturbation with harmful dataset during fine-tuning stage, which recover the safety alignment while maintaining the fine-tuning performance.
+However, as shown in the right of Figure 4, random perturbations significantly impair the fine-tuned model’s performance as the fine-tuning accuracy is also significantly downgraded with the increase of noise intensity.
+[…]
+We need a more carefully crafted post-fine-tuning perturbation.
+As post-fine-tuning perturbation has potential to recover the model to a safe state but it comes with a degradation of model’s generalization performance, we need to explore a better way to craft such perturbation.
+We will discuss this better perturbation crafting method in the following section.
 4 Methodology
-At the fine-tuning stage, we aim to optimize an adaptive perturbation that maximally increases the harmful loss.
+In this section, we discuss our method to craft a better post-fine-tuning perturbation to recover the model from fine-tuning attack.
+To search for such perturbation, we operate an extra optimization at the fine-tuning stage.
+Specifically, at the fine-tuning stage, we aim to optimize an adaptive perturbation that maximally increases the harmful loss.
 This perturbation is then added to the fine-tuned model after the fine-tuning process.
-Therefore, our method is a post-fine-tuning stage solution, formulated as a max-maximize optimization, as follows:
+Formally, our method can be formulated as a max-maximize optimization, as follows:
 $\max_{\boldsymbol{w}}\max_{\boldsymbol{\varepsilon}:\|\boldsymbol{\varepsilon}\|\leq\rho}\lambda(h(\boldsymbol{w}+\boldsymbol{\varepsilon})-h(\boldsymbol{w}))-g(\boldsymbol{w})$ (1)
@@
 where $\eta$ is the learning rate.
-As shown in Algorithm 1, the optimization process consists of four key steps: First, a batch of fine-tuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$ is used to compute the gradient ~$\widetilde{\nabla}g(\boldsymbol{w}_{t})$, where $\widetilde{\nabla}$ notes a batch of gradient.
-Second, a batch of harmful data $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$ is sampled to compute the harmful gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t})$.
-Third, the perturbation is computed (Eq.
-2) and applied to update the harmful gradient, yielding ~$\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$.
-Lastly, the combined gradient ~$\widetilde{\nabla}f(\boldsymbol{w}_{t})$ is computed and used to update the model parameters, with the final perturbation applied to obtain the re-aligned parameters $\boldsymbol{w}_{\varepsilon}$.
 Algorithm 1 Panacea: Adaptive Perturbation Optimization
@@
 1:  for each iteration $t=0,\dots,T-1$ do
-2:    Sample a batch of finetuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
+2:  Sample a batch of fine-tuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
 3:  Sample a batch of harmful data $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
-4:    Compute gradient ~$\widetilde{\nabla}g(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
-5:    Compute gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
+4:  Compute gradient $\widetilde{\nabla}g(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$
+5:  Compute gradient $\widetilde{\nabla}h(\boldsymbol{w}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
 6:  Compute perturbation $\boldsymbol{\varepsilon}_{t}=\rho\frac{\widetilde{\nabla}h(\boldsymbol{w}_{t})}{\|\widetilde{\nabla}h(\boldsymbol{w}_{t})\|}$
-7:    Compute gradient ~$\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
-8:    ~$\widetilde{\nabla}f(\boldsymbol{w}_{t})=\lambda(\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})-\widetilde{\nabla}h(\boldsymbol{w}_{t}))-\widetilde{\nabla}g(\boldsymbol{w}_{t})$
+7:  Compute gradient $\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$ on $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$
+8:  $\widetilde{\nabla}f(\boldsymbol{w}_{t})=\lambda(\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})-\widetilde{\nabla}h(\boldsymbol{w}_{t}))-\widetilde{\nabla}g(\boldsymbol{w}_{t})$
 9:  $\boldsymbol{w}_{t+1}=\boldsymbol{w}_{t}+\eta\widetilde{\nabla}f(\boldsymbol{w}_{t})$
@@
 11:  $\boldsymbol{w}_{\varepsilon}\leftarrow\boldsymbol{w}_{T}+\boldsymbol{\varepsilon}_{T-1}$
-Our proposed algorithm, dubbed as Panacea, is named after an alignment-stage defense Vaccine (Huang et al., 2024e).
+As shown in Algorithm 1, the optimization process consists of four key steps: First, a batch of fine-tuning data $(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$ is used to compute the gradient $\widetilde{\nabla}g(\boldsymbol{w}_{t})$, where $\widetilde{\nabla}$ notes a batch of gradient.
+Second, a batch of harmful data $(\boldsymbol{x}^{\prime}_{t},\boldsymbol{y}^{\prime}_{t})$ is sampled to compute the harmful gradient $\widetilde{\nabla}h(\boldsymbol{w}_{t})$.
+Third, the perturbation is computed (Eq.
+2) and applied to update the harmful gradient, yielding $\widetilde{\nabla}h(\boldsymbol{w}_{t}+\boldsymbol{\varepsilon}_{t})$.
+Lastly, the combined gradient $\widetilde{\nabla}f(\boldsymbol{w}_{t})$ is computed and used to update the model parameters, with the final perturbation applied to obtain the re-aligned parameters $\boldsymbol{w}_{\varepsilon}$.
+Our proposed algorithm, dubbed as Panacea, is named after an alignment-stage defense Vaccine Huang et al.
+(2024e).
 However, we note that these two defenses are fundamentally different.
 Vaccine vaccinates the LLM to enhance its robustness at the alignment stage in order to counter attacks launched after alignment.
-By constrast, our algorithm Panacea belongs to the post-fine-tuning stage, where it introduces an optimized adaptive perturbation to restore the model’s safety alignment after the fine-tuning stage.
-We also show in experiment that Panacea outperforms Vaccine in both the two concerned metrics, i.e., harmful score and fine-tuning accuracy.
-Table 1: Performance comparison of different harm ratio.
-Method Harmful Score($\downarrow$) Finetuning Accuracy($\uparrow$)
-Clean 0.05 0.1 0.15 0.2 0.25 Avg. Clean 0.05 0.1 0.15 0.2 0.25 Avg.
-SFT 5.2 27.5 45.8 56.2 67.0 69.4 45.2 16.4 16.0 16.2 15.4 15.2 16.8 16.0
-Vaccine 2.3 15.6 25.4 40.3 55.2 65.9 34.1 14.2 13.9 13.5 13.0 13.6 13.4 13.6
-RepNoise 2.7 22.2 32.0 42.9 54.0 62.4 36.0 15.7 16.1 15.8 14.9 14.0 14.8 15.2
-Booster 4.6 21.0 42.3 60.7 69.3 72.5 45.1 17.8 17.8 17.6 17.1 16.2 14.7 16.9
-Panacea 1.8 9.9 20.1 29.1 34.8 46.3 23.7 15.0 16.3 16.7 17.0 16.2 17.4 16.3
+By contrast, our algorithm Panacea belongs to the post-fine-tuning stage, where it introduces an optimized adaptive perturbation to restore the model’s safety alignment after the fine-tuning stage.
+Since Panacea does not require access to the alignment stage, it can be directly applied to already aligned LLMs such as Llama2-7B-Chat.
+We present experimental results on such aligned models in Table 4, demonstrating the broad applicability of our method.
+Furthermore, our experiments show that Panacea outperforms Vaccine on both key metrics—harmful score and fine-tuning accuracy.
+Of note, when we prepare the camera ready of this paper, we find a previous work Security Vectors  Zhou et al.
+(2023) follow a similar idea with Panacea.
+Both Security Vectors and Panacea aim to make sure that the harmful loss can be sufficiently reduced during fine-tuning, and then add a perturbation after fine-tuning to remove the harmful knowledge.
+However, there is two difference between Security vector and Panacea.
+The first implementation difference is that security vector learns the perturbation before fine-tuning while Panacea learns the perturbation during fine-tuning.
+The second difference, which is the major difference is what the perturbation is and how it leads to increase of harmful loss.
+Panacea aim to find a perturbation that maximally increase the loss harmful loss and thereby adding the perturbation can sufficiently unlearn harmful knowledge.
+In contrast, the goal of security vector is to distill harmful knowledge into a harmful component before fine-tuning, and removing this component (which can be seen as a perturbation as well) to increase harmful loss.
+However, their formulation can not explicitly guarantee that the increase of harmful loss is maximized after adding perturbation (i.e., de-activating their security vector).
 5 Experiment
@@
 Three distinct datasets are utilized: the alignment dataset, harmful dataset, and fine-tuning dataset.
-The alignment dataset and harmful dataset are derived from the RepNoise (Rosati et al., 2024b), which extracts subsets from the BeaverTails dataset (Ji et al., 2023).
+The alignment dataset and harmful dataset are derived from the RepNoise Rosati et al.
+(2024b), which extracts subsets from the BeaverTails dataset Ji et al.
+(2023).
 Specifically, 5,000 examples are sampled for the alignment dataset, and 1,000 examples for the harmful dataset.
-The fine-tuning dataset is constructed from four downstream fine-tuning tasks: GSM8K (Cobbe et al., 2021), SST2 (Socher et al., 2013), AlpacaEval (Li et al., 2023), and AGNEWS (Zhang et al., 2015), with 1,000 samples collected from each task (700 samples from AlpacaEval).
-To simulate the harmful fine-tuning attack, we combine $p$ (percentage) of harmful data with $(1-p)$ of benign fine-tuning data, and $p$ is set to 0.1 by default.
-The harmful data is also sourced from BeaverTails (Ji et al., 2023) and does not overlap with the harmful dataset.
+The fine-tuning dataset is constructed from four downstream fine-tuning tasks: GSM8K Cobbe et al.
+(2021), SST2 Socher et al.
+(2013), AlpacaEval Li et al.
+(2023), and AGNEWS Zhang et al.
+(2015), with 1,000 samples collected from each task (700 samples from AlpacaEval).
+To simulate the harmful fine-tuning attack, we combine $p$ (percentage) of harmful data with (1p)(1-p) of benign fine-tuning data, and $p$ is set to 0.1 by default.
+The harmful data is also sourced from BeaverTails Ji et al.
+(2023) and does not overlap with the harmful dataset.
 Baseline.
@@
 The SFT method is the vanilla supervised training using the alignment dataset.
-Vaccine (Huang et al., 2024e) applies supervised training on the alignment dataset while introducing additional perturbations.
-Both RepNoise (Rosati et al., 2024b) and Booster (Huang et al., 2024b) utilize the alignment and harmful datasets for supervised and adversarial training.
+Vaccine Huang et al.
+(2024e) applies supervised training on the alignment dataset while introducing additional perturbations.
+Both RepNoise Rosati et al.
+(2024b) and Booster Huang et al.
+(2024b) utilize the alignment and harmful datasets for supervised and adversarial training.
 All four baseline methods are trained exclusively during the alignment stage.
+Antidote Huang et al.
+(2024a)is a post-fine-tuning stage baseline that utilizes a harmful dataset after the fine-tuning stage.
 Our proposed method incorporates training with the harmful dataset during the fine-tuning stage after vanilla alignment training.
-More details are in Appendix C.
-Metric. Following  (Huang et al., 2024e), two metrics are used to evaluate the model’s performance.
+More details are in Appendix E.
+Metric. Following  Huang et al. (2024e), two metrics are used to evaluate the model’s performance.
   • HS (Harmful Score): It reflects the frequency with which the model generates harmful content when handling malicious instructions.
-Harmful Score is determined by a moderation model, provided by  (Ji et al., 2023), which assesses whether the model’s output is harmful in response to a given harmful instruction.
-And a sample of 1,000 instructions is drawn from the BeaverTails (Ji et al., 2023) test set to compute this metric.
-  • FA (Finetuning Accuracy): It refers to the accuracy on various downstream fine-tuning tasks.
+Harmful Score is determined by a moderation model, provided by  Ji et al.
+(2023), which assesses whether the model’s output is harmful in response to a given harmful instruction.
+And a sample of 1,000 instructions is drawn from the BeaverTails Ji et al.
+(2023) test set to compute this metric.
+  • FA (Fine-tuning Accuracy): It refers to the accuracy on various downstream fine-tuning tasks.
 Samples are sampled from the test sets of GSM8K, SST2, AlpacaEval, and AGNEWS, with 1,000, 872, 1,000, and 105 samples, respectively, used to compute this metric.
-More details are in Appendix B.
+Details are in Appendeix D.1
 Implementation Details.
-For efficient training, the approach follows the methodology (Huang et al., 2024e), utilizing LoRA (Hu et al., 2021) with a rank of 32 and an alpha value of 4.
-And the optimizer is AdamW (Loshchilov et al., 2017).
-During the alignment stage, the learning rate is set to $45e-$ 4, the batch size is 10, and the training is performed for 20 epochs.
-For the fine-tuning stage, the learning rate is set to $52e-$ 5, with a batch size for 10 and a training epoch for 20.
-These settings are applied uniformly across all datasets and baselines, with the default dataset being GSM8K (Cobbe et al., 2021) and the default model being Llama2-7B (Touvron et al., 2023) following (Rosati et al., 2024b; Huang et al., 2024b).
-To verify the robustness of the approach, two state-of-the-art LLMs, Gemma2-9B (Team et al., 2024) and Qwen2-7B (Yang et al., 2024), are included in the evaluation.
-Table 2: Performance comparison of different fine-tuning tasks.
-GSM8K SST2 AlpacaEval AGNEWS Average
-Method HS FA HS FA HS FA HS FA HS FA
-SFT 45.8 16.2 55.5 94.04 23.1 46.15 54.3 83.5 44.68 59.97
-Vaccine 25.4 13.5 53.8 93.35 34.7 37.50 54.9 85.1 42.20 57.36
-RepNoise 32.0 15.8 61.5 93.81 24.4 44.23 58.1 85.0 44.00 59.71
-Booster 42.3 17.6 49.5 93.23 21.9 45.19 46.5 85.4 40.05 60.36
-Panacea 20.1 16.7 32.2 92.78 18.7 48.08 44.5 81.1 28.88 59.66
+For efficient training, the approach follows the methodology Huang et al.
+(2024e), utilizing LoRA Hu et al.
+(2021) with a rank of 32 and an alpha value of 4.
+And the optimizer is AdamW Loshchilov et al.
+(2017).
+During the alignment stage, the learning rate is set to 5e45e-4, the batch size is 10, and the training is performed for 20 epochs.
+For the fine-tuning stage, the learning rate is set to 2e52e-5, with a batch size for 10 and a training epoch for 20.
+These settings are applied uniformly across all datasets and baselines, with the default dataset being GSM8K Cobbe et al.
+(2021) and the default model being Llama2-7B Touvron et al.
+(2023) following Rosati et al.
+(2024b); Huang et al.
+(2024b).
+To verify the robustness of the approach, two state-of-the-art LLMs, Gemma2-9B Team et al.
+(2024) and Qwen2-7B Yang et al.
+(2024a), are included in the evaluation.
 5.2 Main Results
-In this sub-section, we conduct a comprehensive evaluation of Panacea for the effectiveness and generalization.
+We conduct a comprehensive evaluation of Panacea for the effectiveness and generalization.
 Harmful Ratio.
-Fine-tuning datasets with different harmful ratios are employed, specifically 0 (clean), 0.05, 0.1, 0.15, 0.2, and 0.25.
+Fine-tuning datasets with different harmful ratios are employed, specifically 0 (Clean), 0.05, 0.1, 0.15, 0.2.
 The results are presented in Table 1, Panacea achieves the lowest harmful score across different harmful ratios while maintaining competitive fine-tuning performance (ranked as the second-best on average), indicating that the expected adaptive perturbation is obtained, and the analysis is in Sec 5.3.
-Compared to SFT method, it reduces the harmful score by an average of 21.5% and improves fine-tuning accuracy by 0.3%.
+Compared to SFT method, it reduces the harmful score by an average of 21.2% and improves fine-tuning accuracy by 0.4%.
 Furthermore, as the harmful ratio increases, Panacea consistently maintains a lower harmful score compared to other methods.
@@
 However, since Panacea is designed to weaken harmful loss, its fine-tuning performance on clean data (without explicit harmful loss) is slightly reduced, while Panacea still achieves the lowest harmful score on this benign fine-tuing.
+Table 1: Performance comparison of different harm ratio.
+Method Harmful Score($\downarrow$) Fine-tuning Accuracy($\uparrow$)
+Clean 0.05 0.1 0.15 0.2 Avg. Clean 0.05 0.1 0.15 0.2 Avg.
+SFT 5.2 27.5 45.8 56.2 67.0 40.3 16.4 16.0 16.2 15.4 15.2 15.8
+Vaccine 2.3 15.6 25.4 40.3 55.2 27.8 14.2 13.9 13.5 13.0 13.6 13.6
+RepNoise 2.7 22.2 32.0 42.9 54.0 30.8 15.7 16.1 15.8 14.9 14.0 15.3
+Booster 4.6 21.0 42.3 60.7 69.3 39.6 17.8 17.8 17.6 17.1 16.2 17.3
+Antidote 2.3 14.0 27.2 32.7 35.3 22.3 17.9 16.2 15.3 16.3 16.2 16.3
+Panacea 1.8 9.9 20.1 29.1 34.8 19.1 15.0 16.3 16.7 17.0 16.2 16.2
 Fine-tuning Tasks.
@@
 Overall, Panacea exhibits strong generalization across different fine-tuning tasks.
+Table 2: Performance comparison of different fine-tuning tasks.
+GSM8K SST2 AlpacaEval AGNEWS Average
+Method HS FA HS FA HS FA HS FA HS FA
+SFT 45.8 16.2 55.5 94.04 23.1 46.15 54.3 83.5 44.7 59.9
+Vaccine 25.4 13.5 53.8 93.35 34.7 37.50 54.9 85.1 42.2 57.3
+RepNoise 32.0 15.8 61.5 93.81 24.4 44.23 58.1 85.0 44.0 59.7
+Booster 42.3 17.6 49.5 93.23 21.9 45.19 46.5 85.4 40.0 60.3
+Antidote 27.2 15.3 43.5 93.58 19.4 33.01 47.5 84.9 34.4 56.6
+Panacea 20.1 16.7 32.2 92.78 18.7 48.08 44.5 81.1 28.9 59.6
 Mainstream LLMs.
@@
 Booster 42.3 17.6 26.4 52.6 13.5 65.6 27.4 45.3
+Antidote 27.2 15.3 22.2 61.0 9.6 65.2 19.6 47.1
 Panacea 20.1 16.7 10.7 52.1 5.2 65.4 12.0 44.7
+Aligned LLMs.
+Since our method operates at the post-fine-tuning stage, it can be directly applied to already safety-aligned large language models (LLMs), such as Llama2-7B-Chat, Gemma2-9B-It, and Qwen2-7B-Instruct.
+In contrast, methods like Vaccine must be applied during the alignment stage and are therefore inapplicable to pre-aligned LLMs.
+We compare our approach with two other post-fine-tuning methods: Safe LoRA Hsu et al.
+(2024) and Antidote Huang et al.
+(2024a).
+As shown in Table 4, Panacea reduces the average harmful score by about 20% compared to the SFT baseline, while incurring only a 0.2% drop in accuracy.
+Moreover, compared to other post-fine-tuning stage methods, our method consistently achieves greater reductions in harmful scores.
+On the Gemma model, Panacea achieves nearly a 15% lower harmful score than the best competing method, demonstrating the effectiveness of the optimized perturbation introduced by our approach.
+Table 4: Performance comparison on aligned LLMs.
+Llama2-7B-Chat Gemma2-9B-It Qwen2-7B-Instruct Average
+Method HS FA HS FA HS FA HS FA
+SFT 47.2 20.0 54.4 77.0 28.9 67.2 43.5 54.7
+Safe LoRA 46.8 20.5 56.8 76.3 29.9 67.5 44.5 54.8
+Antidote 37.3 20.4 31.0 76.9 23.3 59.3 30.5 52.2
+Panacea 35.7 17.2 15.6 80.3 19.8 66.1 23.7 54.5
+Evaluation Benchmark.
+To better validate the effectiveness of Panacea, we conduct additional evaluations on both Sorry-Bench Xie et al.
+(2025) and the AdvBench Zou et al.
+(2023) under different harmful ratios using LLaMA-2-7B.
+We also include the ConstrainSFT method that is proposed in Qi et al.
+(2024a) Section 4.1.
+The evaluation results are presented in Table 5.
+For Sorry-Bench, Fulfillment Rate (FR) is used as the metric (lower is better ↓).
+As the results show, Panacea consistently achieves the best performance across all three settings, demonstrating its effectiveness and generalizability.
+In particular, on AdvBench, the harmful score remains as low as 10.58% even under the most extreme setting.
+Table 5: Evaluations on AdvBench and Sorry-Bench under diffusion harmful ratios.
+AdvBench HS (ratio=0.05) HS (ratio=0.1) HS (ratio=0.15) HS (ratio=0.2)
+SFT 7.50 17.69 37.50 48.65
+Vaccine 25.19 49.62 59.23 71.35
+RepNoise 3.46 8.46 20.38 40.00
+ConstrainSFT 4.04 12.12 20.58 34.81
+Panacea 0.00 1.54 5.19 10.58
+Sorry-Bench FR (ratio=0.05) FR (ratio=0.1) FR (ratio=0.15) FR (ratio=0.2)
+SFT 45.23 57.50 65.00 70.23
+Vaccine 34.32 49.77 53.63 65.45
+RepNoise 60.91 66.59 75.23 81.14
+ConstrainSFT 45.23 59.55 61.36 66.36
+Panacea 33.41 42.05 49.55 54.32
+Harmful Data from Different Sources.
+We conducted the experiment that harmful data is from different sources.
+Specifically, the harmful data used during the defense phase remains from BeaverTails Ji et al.
+(2023), while the harmful data used for fine-tuning in the attack phase is replaced with data from LLM-LAT Sheshadri et al.
+(2024).
+And the harmful score is evaluated using test set from AdvBench Zou et al.
+(2023).
+The experimental results are shown in Table 6.
+As shown, Panacea significantly reduces the harmful score compared to other methods, even when the harmful data come from different sources.
+This result further demonstrates the effectiveness of our method.
+Table 6: Performance comparison under different harmful data sources.
+Method HS (ratio=0.05) HS (ratio=0.1) FA (ratio=0.05) FA (ratio=0.1)
+SFT 74.62 84.81 16.5 15.9
+Vaccine 48.65 73.65 15.2 13.5
+RepNoise 61.73 81.54 15.4 14.7
+ConstrainSFT 60.00 86.35 15.0 15.5
+Panacea 11.73 41.73 17.1 17.1
 5.3 Perturbation Analysis
@@
 For the complicated AlpacaEval dataset, it even improves fine-tuning performance by 3.85%, while only Qwen2-7B shows a slight decrease of 0.3% in other LLM experiments.
-Our post-fine-tuning perturbation successfully restores the model’s safety alignment without negatively affecting fine-tuning performance, and in some cases, it even enhances model performance as reducing harmful behaviors.
-Table 4: Ablation study on different fine-tuning tasks. “w/o” denotes that the adaptive perturbation obtained is not applied after the training process.
+Table 7: Ablation study on different fine-tuning tasks. “w/o" denotes that the adaptive perturbation obtained is not applied after the training.
 GSM8K SST2 AlpacaEval AGNEWS
@@
 Panacea 20.1 16.7 32.2 92.78 18.7 48.08 44.5 81.1
-Table 5: Ablation study on different LLMs. ”w/o” denotes that the adaptive perturbation obtained is not applied after the training process.
+Table 8: Ablation study on different LLMs. "w/o" denotes that the adaptive perturbation obtained is not applied after the training.
 Llama2-7B Gemma2-9B Qwen2-7B
@@
 Panacea 20.1 16.7 10.7 52.1 5.2 65.4
+Our perturbation successfully restores the model’s safety alignment without negatively affecting fine-tuning performance, and it even enhances model performance as reducing harmful behaviors.
+Layer-wise safety property.
+Visualization analysis is conducted by visualizing the weights of the perturbations obtained in Table 8.
+We sum the absolute values of the parameters at each layer, to compute the magnitude of changes by perturbation to the original model.
 Figure 5: Parameter weights of different LLMs.
-The parameters in the earlier layers of Llama2-7B (blue) have larger weights, while Gemma2-9B (yellow) and Qwen (purple) have larger weights in the middle and later layers.
-Visualization.
-Visualization analysis is conducted by visualizing the weights of the perturbations obtained in Table 5.
-The result in Figure 5 reveals the layers of different LLMs that are most critical for model’s safety.
-Specifically, in Llama2-7B, the adaptive perturbation has a larger weight in the earlier layers and a smaller effect on the later layers, suggesting that the earlier layers are more critical for the model’s safety.
-This observation aligns with previous research (Du et al., 2024a; Rosati et al., 2024b; Liu et al., 2024a; Yi et al., 2024b).
+The parameters in the earlier layers of Llama2-7B (blue) have larger weights, while Gemma2-9B (yellow) and Qwen2-7B (purple) have larger weights in the middle and later layers.
+The result in Figure 5 reveals different layers exhibit different affinity towards safety task.
+In Llama2-7B, the adaptive perturbation has a larger weight in the earlier layers and a smaller effect on the later layers, suggesting the earlier layers are more critical for the model’s safety.
+This observation aligns with multiple previous research Du et al.
+(2024a); Rosati et al.
+(2024b); Liu et al.
+(2024a); Yi et al.
+(2024b); Li et al.
+(2024a); Zhao et al.
+(2025b) (See Appendix C for detailed discussions).
 Additionally, we observe that in Gemma2-9B, the middle and final layers hold greater importance for safety, while in Qwen2-7B, the safety importance gradually increases across layers, reaching its peak in the final layers.
 This may indicate that the model implements stricter safety measures in the output layer.
-Therefore, our experiments also suggest that these models may require targeted defense mechanisms for specific layers or precautions during fine-tuning to prevent the disruption of layers that are crucial for safety.
-Figure 6: Model statistics (Top: harmful score, Middle: harmful training loss, Bottom: harmful testing loss) after fine-tuning on fine-tuning dataset(10% of data is harmful) for different steps.
+Therefore, our experiments also suggest these models may require targeted defense for specific layers or precautions during fine-tuning to prevent the disruption of layers that are crucial for safety.
 5.4 Statistical Analysis
@@
 Thus, during the evaluation phase, Panacea adds the perturbation optimized at specific step and subtracts it after the evaluation is finished.
+Figure 6: Model statistics (Top: harmful score, Middle: harmful training loss, Bottom: harmful testing loss) after fine-tuning on fine-tuning dataset(10% of data is harmful) for different steps.
 Harmful Score.
-As shown in the top figure, both Panacea and SFT exhibit an increase in harmful scores during the fine-tuning process.
+In the top figure, both Panacea and SFT exhibit an increase in harmful scores during the fine-tuning process.
 However, the harmful score increase for Panacea is generally smaller.
@@
 Harmful Training Loss.
-As shown in the middle figure, since Panacea does not apply additional defense during the alignment stage, it starts with the same harmful training loss as SFT.
-Although both harmful training losses decrease, Panacea exhibits a smaller reduction, reaching a higher harmful training loss than SFT in the end.
-With the relatively high loss, Panacea’s harmful score shows a significant improvement.
+In the middle figure, since Panacea does not apply additional defense during the alignment stage, it starts with the same harmful training loss as SFT.
+Panacea exhibits a smaller reduction, reaching a higher harmful training loss than SFT in the end.
+With the high loss, Panacea’s harmful score shows a significant improvement.
 Finally, we observe an interesting phenomenon that the harmful training loss and harmful score of Panacea follow the same trend in the last few hundred steps.
 We attribute this to the near-successful optimization of the adaptive perturbation at this stage.
-See more analysis of harmful loss in Appendix B.3.
+See more analysis of harmful loss in Appendix D.3.
 Harmful Testing Loss.
-As shown in the bottom figure, the model’s harmful testing loss exhibits a similar trend to the harmful training loss.
+In the bottom figure, the model’s harmful testing loss exhibits a similar trend to the harmful training loss.
 This is because the data used during training and testing come from the same distribution, and the methods applied during training are transferable to unseen data.
-5.5 System Evaluation
-Table 6 presents a comparison of the clock time and GPU memory usage during training for different methods.
-Clock Time.
-Panacea requires only 0.25 more hours than SFT for total training time, and it outperforms other state-of-the-art methods in terms of time efficiency.
-Specifically, other methods double or even more than double the time spent on the time-consuming alignment stage.
-We acknowledge that the adaptive perturbation optimized during the fine-tuning stage in Panacea significantly improves safety, but it incurs an additional time cost—more than doubling the time.
-This is because we perform two additional gradient computations during training.
-However, compared to alignment-stage methods, our approach offers a time advantage as it can be directly applied to the aligned model, making it more practical even time-efficient.
-GPU Memory.
-Panacea achieves the lowest memory usage.
-In contrast, Vaccine and RepNoise introduce an additional 9.52GB and 37.57GB of memory usage compared to SFT.
-Panacea reduces memory usage slightly compared to SFT, which we believe is due to the reduced gradient size after the final gradient summation.
-Table 6: Comparison of methods in terms of clock time and GPU memory. The second best results are underlined. ”Align.” and ”Fine.” suggest alignment and fine-tuning.
-Methods Clock Time (Hour) GPU Memory (GB)
-Align. Fine. Sum Align. Fine. Max
-SFT 0.58 0.17 0.75 34.90 32.90 34.90
-Vaccine 1.12 0.17 1.29 44.42 32.90 44.42
-Repnoise 2.67 0.17 2.84 72.47 32.90 72.47
-Booster 1.87 0.17 2.04 35.98 32.90 35.98
-Panacea 0.58 0.42 1.00 34.90 32.86 34.90
 5.5 Hyper-parameter Analysis
@@
 This is in expectation, because with a larger $\lambda$, we put more weights in maximize the harmful loss after adding perturbation, and therefore resulting in a lower harmful score.
-On the other hand, we do observe that when $\lambda$ is too large, the fine-tune accuracy will degrade.
-Therefore, a validation dataset is required to select a proper $\lambda$ based on the need of the deployment.
-Table 7: Impact of perturbation intensity $\rho$ over Panacea.
-As $\rho$ increases, the harmful score decreases.
-However, when $\rho$ continues to increase, the harmful score decreases further, but the fine-tuning accuracy also begins to decrease.
+Besides, we observe that when $\lambda$ is too large, the fine-tune accuracy will degrade.
+Table 9: Impact of perturbation intensity $\rho$.
 $\rho=0$ $\rho=0.1$ $\rho=0.5$ $\rho=1$ $\rho=2$ $\rho=5$
@@
 FA 16.2 16.3 18.0 16.7 13.5 6.0
-Table 8: Impact of regularizer’s intensity $\lambda$ over Panacea. As $\lambda$ increases, the harmful score and the fine-tuning accuracy decrease.
+Table 10: Impact of regularizer’s intensity $\lambda$.
 $\lambda=0$ $\lambda=1\times 10^{-4}$ $\lambda=0.001$ $\lambda=0.01$
@@
 The comprehensive experiments demonstrate the effectiveness and generalization of Panacea and the visualization of the adaptive perturbation reveals the different lays in various LLMs have distinct safety coefficients.
-Impact Statement
-This paper presents work whose goal is to address the harmful fine-tuning and make LLMs helpful and harmless.
-We acknowledge that the phenomena or issues identified in this paper may pose potential risks.
-Disclaimer: this paper contains red-teaming data (from open dataset) and model-generated content that can be offensive in nature.
+Acknowledgments and Disclosure of Funding
+Li Shen is supported by National Key R&D Projects (NO.
+2024YFC3307100), NSFC Grant (No.
+62576364), Shenzhen Basic Research Project (Natural Science Foundation) Basic Research Key Project (NO.
+JCYJ20241202124430041), CCF-DiDi GAIA Collaborative Research Funds (NO.
+CCF-DiDi GAIA 202419 and CCF-DiDi GAIA 202519)
 Appendix A Proof of Inner Optimization
@@
 This completes the proof.
-Appendix B More Detail of Experiments.
-B.1 Experiment Details
+Besides, it is necessary to impose magnitude constraints on $\boldsymbol{\varepsilon}$: Without this constraint, the optimization objective $\boldsymbol{\varepsilon}^{T}\nabla h(\boldsymbol{w})$ becomes unbounded, as $\boldsymbol{\varepsilon}$ can be scaled arbitrarily in the direction of $\nabla h(\boldsymbol{w})$, leading to an infinite increase in the objective value.
+In such a case, the optimization has no finite optimal solution.
+Therefore, the L2 norm constraint is not just a technical detail—it is necessary to ensure the existence of a valid and bounded solution.
+Appendix B More Analysis.
+Model Size.
+LLaMA2 is available in 7B, 13B, and 32B versions, while Qwen2 is available in 0.5B, 1.5B, and 72B.
+Therefore, we conducted experiments using the larger LLaMA2-13B and the smaller Qwen2-1.5B to ensure a more comprehensive evaluation.
+As shown in Table 11, our method consistently reduces the harmful score across LLMs of different sizes, achieving an average reduction of 11.8%, while also attaining the highest Fine-tuning Accuracy.
+Table 11: Performance comparison of different sizes.
+Llama2-13B Qwen2-1.5B Average
+Method HS FA HS FA HS FA
+SFT 36.2 39.2 46.9 25.3 41.6 32.3
+Vaccine 22.4 32.3 35.5 17.6 28.9 25.0
+RepNoise 31.2 35.7 36.6 25.7 33.9 30.7
+Panacea 17.8 40.9 26.4 25.2 22.1 33.1
+Harmful Ratio.
+Building on Table 1, we further increase the ratio of harmful data in the fine-tuning dataset.
+As shown in Table 12, Panacea consistently achieves the lowest harmful score and better fine-tuning accuracy compared to other baselines.
+Panacea also shows a performance drop at a ratio of 0.5 (while still outperforming other baselines).
+Nevertheless, such high-ratio settings are not the focus of our study, as we primarily investigate scenarios where the dataset contains only a small fraction of harmful data.
+Table 12: Performance comparison on higher ratio.
+0.3 0.4 0.5
+Method HS FA HS FA HS FA
+Vaccine 70.6 13.0 72.6 11.0 76.0 11.9
+RepNoise 66.7 14.5 70.5 14.3 73.4 12.7
+Panacea 49.4 16.0 61.7 16.7 65.5 15.0
+Optimization Objectives.
+For Panacea, the optimization objective is to maximize $\lambda\left(h(\boldsymbol{w}+\boldsymbol{\varepsilon})-h(\boldsymbol{w})\right)-g(\boldsymbol{w})$.
+We conduct experiments under the same setup using the objective function $\max\lambda h(\boldsymbol{w})-g(\boldsymbol{w})$, where the only hyperparameter is $\lambda$ to trade off the two loss terms.
+The results are shown in Table 13.
+Table 13: Results under different $\lambda$ for the objective $\lambda h(\boldsymbol{w})-g(\boldsymbol{w})$ and Panacea.
+$\lambda h(\boldsymbol{w})-g(\boldsymbol{w})$ $\lambda=0.0001$ $\lambda=0.001$ $\lambda=0.01$ $\lambda=0.1$ $\lambda=1$ Panacea
+HS $\downarrow$ 45.4 44.6 39.1 fail fail 20.1
+FA $\uparrow$ 16.3 16.3 16.5 11.0 7.9 16.7
+We can observe that the best harmful score achieved by optimizing $\lambda h(\boldsymbol{w})-g(\boldsymbol{w})$ is 39.1, while Panacea achieves a much lower score of 20.1.
+We believe this is due to the inherent difficulty of optimizing two opposing objectives (i.e., increasing $h(\boldsymbol{w})$ while decreasing $g(\boldsymbol{w})$ using a single shared parameter $\boldsymbol{w}$).
+In contrast, Panacea performs gradient ascent primarily through adaptive perturbations, which allows it to better reduce harmfulness without degrading utility.
+As $\lambda$ increases, directly maximizing $h(\boldsymbol{w})$ degrades the model, where the model produces no meaningful output in “fail” cases.
+However, Panacea, as shown in Table 10, does not suffer from such failure.
+This is because in Panacea’s formulation (Eq.
+1), the term $h(\boldsymbol{w})$ acts as a regularization that prevents excessive optimization, and the inner perturbation is further constrained by a fixed L2 norm bound $\rho$.
+System Evaluation Table 14 presents a comparison of the clock time and GPU memory usage on A100-80GB during training for different methods.
+Table 14: Comparison of methods in terms of clock time and GPU memory. The second best results are underlined. "Align." and "Fine." suggest alignment and fine-tuning.
+Methods Clock Time (Hour) GPU Memory (GB)
+Align. Fine. Sum Align. Fine. Max
+SFT 0.58 0.17 0.75 34.90 32.90 34.90
+Vaccine 1.12 0.17 1.29 44.42 32.90 44.42
+Repnoise 2.67 0.17 2.84 72.47 32.90 72.47
+Booster 1.87 0.17 2.04 35.98 32.90 35.98
+Panacea 0.58 0.42 1.00 34.90 32.86 34.90
+Clock Time.
+Panacea requires only 0.25 more hours than SFT for total training time, and it outperforms other state-of-the-art methods in terms of time efficiency.
+Specifically, other methods double or even more than double the time spent on the time-consuming alignment stage.
+We acknowledge that the adaptive perturbation optimized during the fine-tuning stage in Panacea significantly improves safety, but it incurs an additional time cost—more than doubling the time.
+This is because we perform two additional gradient computations during training.
+However, compared to alignment-stage methods, our approach offers a time advantage as it can be directly applied to the aligned model, making it more practical even time-efficient.
+GPU Memory.
+Panacea achieves the lowest memory usage.
+In contrast, Vaccine and RepNoise introduce an additional 9.52GB and 37.57GB of memory usage compared to SFT.
+Panacea reduces memory usage slightly compared to SFT, which we believe is due to the reduced gradient size after the final gradient summation.
+Extra Computes Required.
+As shown in Algorithm 1, our post-fine-tuning perturbation is actually computed during the fine-tuning stage and simply applied to the model parameters at the post-fine-tuning stage.
+In our setting, the user fine-tuning runs for 20 epochs, and the perturbation optimization is performed concurrently within the same epochs.
+Therefore, the extra computation does not exceed the normal user fine-tuning budget.
+We acknowledge the introduction of additional computation.
+We evaluated two newly proposed baselines Qi et al.
+(2024a); Li et al.
+(2025c) (the results are shown below), and also measured their additional computation time.
+For ConstrainSFT Qi et al.
+(2024a), the defense introduces 0.09 hours of runtime, but since it requires an aligned model as a reference model, the memory consumption reaches 48.2 GB, which is 15.3 GB more than Panacea.
+Moreover, its reduction in harmful score is less significant than Panacea.
+For SaLoRA Li et al.
+(2025c), the preprocessing step of setting the weights of the safety module introduces an additional 0.33 hours, which is 0.08 hours more than Panacea.
+Table 15: Comparison of harmful scores and finetune accuracy under different harmful ratios.
+Method HS (ratio=0.05) HS (ratio=0.1) HS (ratio=0.15) HS (ratio=0.2) FA (ratio=0.05) FA (ratio=0.1) FA (ratio=0.15) FA (ratio=0.2)
+ConstrainSFT 21.0 35.2 48.1 57.3 15.5 15.2 16.8 14.2
+SaLoRA 15.9 1.1 1.1 0.9 1.4 3.9 2.5 2.4
+Panacea 9.9 20.1 29.1 34.8 16.3 16.7 17.0 16.2
+Table 16: Comparison of extra clock time under different harmful ratios.
+Method Extra Clock Time (h) GPU Memory (GB)
+ConstrainSFT 0.09 48.20
+SaLoRA 0.33 32.90
+Panacea 0.25 32.86
+Perturbation of Different Scales.
+In Table 8 and Table 8, we conduct an ablation study on whether to apply the post-fine-tuning perturbation obtained through optimization.
+Here, we further analyze the results of incorporating perturbations with different scales in Table 17.
+Table 17: Performance under different perturbation scales.
+Scale 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0 1.2 1.5 2.0 5.0
+HS $\downarrow$ 42.5 40.2 38.9 36.0 32.9 30.1 27.4 26.1 22.8 20.1 17.3 12.9 8.3 2.1
+FA $\uparrow$ 16.0 17.0 17.2 17.3 17.8 17.8 17.8 17.8 17.3 16.7 16.9 16.8 15.1 6.3
+We can observe from the results that as the scale of the added perturbation increases, the model’s HS score decreases accordingly.
+At the same time, different perturbation scales have minimal impact on FA, except when the scale is extremely large (e.g., 5.0), which slightly affects model performance.
+This further demonstrates that the perturbation obtained through our optimization is close to optimal.
+Comparision on Model Weights. Compared to pre-trained model, which contains 6,738M parameters, Panacea only introduces 25M parameters, accounting for just 0.37% of the model weights.
+Topic Distribution.
+Sorry-Bench consists of 44 topics, each containing 10 samples, resulting in a total of 440 samples in the dataset.
+As shown in the Table 18, all methods produced harmful responses in Topic 31 (Military Use).
+Additionally, Topic 29 (False Advertising) triggered harmful responses in all 10 samples for the RepNoise, ConstrainSFT, and Panacea methods.
+Therefore, defenses should pay particular attention to these two topics.
+Table 18: Topic-level analysis on Sorry-Bench.
+Sorry-Bench Top 5 Topic ids Top 5 Topic Violation Counts
+SFT 31, 32, 34, 42, 26 10, 10, 10, 10, 9
+Vaccine 4, 31, 11, 20, 12 9, 9, 8, 8, 7
+RepNoise 29, 31, 34, 41, 42 10, 10, 10, 10, 10
+ConstrainSFT 27, 29, 25, 26, 31 10, 10, 9, 9, 9
+Panacea 29, 16, 31, 34, 41 10, 8, 8, 8, 8
+We divided the responses in AdvBench into 14 topics in total.
+As shown in the Table 19, all methods show the highest number of harmful responses in the topic “violence, aiding_and_abetting, incitement”, suggesting they deserve special attention in defense design.
+Compared to other methods, Panacea significantly reduces the number of harmful responses across all major topics, especially in the top-1 topic where it drops from 298 (Vaccine) to just 45, demonstrating strong safety recovery capability.
+Table 19: Topic-level analysis on AdvBench.
+AdvBench Top1 Topic Violation Count Top2 Topic Violation Count Top3 Topic Violation Count
+SFT violence,… 215 financial_crime,… 140 unethical_behavior 35
+Vaccine violence,… 298 financial_crime,… 152 unethical_behavior 55
+RepNoise violence,… 175 financial_crime,… 108 unethical_behavior 22
+ConstrainSFT violence,… 149 financial_crime,… 110 unethical_behavior 18
+Panacea violence,… 45 financial_crime,… 25 unethical_behavior 7
+Appendix C Perturbation of Different Layers
+As shown in Figure 5, Panacea demonstrates that for the Llama2-7B model on the GSM8K task, the early and middle layers contribute significantly to model safety.
+This observation aligns with the findings of several previous study (though these previous study use different statistical method to demonstrate the safety criticalness of each layer:
+NLSR Yi et al. (2024b). This method identifies that targeting safety-critical neurons within layers 8 to 11 results in the most substantial reduction in the harmful score on GSM8K.
+RepNoise Rosati et al.
+(2024b).
+RepNoise measures the safety relevance of different layers by training a linear probe on the activations at each layer to predict whether an answer is harmful.
+The results show that the middle layers (around layer 10) achieve the highest probe accuracy, indicating they contain the most information about harmfulness.
+This is consistent with our observation that the middle layers play a more critical role in ensuring model safety.
+Targeted Vaccine Liu et al.
+(2024a).
+Targeted Vaccine assesses the safety importance of different layers by adding perturbations to various subsets of layers and measuring the resulting harmful score.
+The results show that applying perturbations to the early and middle layers (around the first 20 layers) yields the best defense performance, while including all layers—especially the last few—degrades effectiveness.
+This aligns with our observation that the early to middle layers play the most critical role in ensuring model safety.
+SPPFT  Li et al.
+(2024a).
+SPPFT compares the pre-trained and aligned versions of Llama models and finds that while the pre-trained models show no noticeable difference between N-N and N-M pairs across all layers, the aligned models exhibit a clear divergence in the middle layers (around layers 10–20).
+This suggests that these middle layers are where the model begins to differentiate between normal and malicious queries, highlighting their critical role in achieving safety alignment.
+This observation is consistent with our finding that safety-relevant signals primarily emerge in the middle of the layers.
+One difference from our findings is that they observe the later layers to be more important than the earlier ones.
+SWAT  Du et al.
+(2024a).
+SWAT assesses the safety importance of each layer by perturbing specific modules (e.g., Q/K/V) at different layers and measuring the resulting performance drop.
+The results show that perturbations in the early to middle layers (e.g., layers 0–12) cause the most significant degradation, which aligns with our observation that the early and middle layers are more critical for model safety.
+RSN-Tune  Zhao et al.
+(2025b).
+RSN-Tune evaluates the safety importance of different layers by progressively deactivating safety neurons and measuring changes in the attack success rate (ASR).
+The results show that disabling the first 10 layers of LLama2-7B-Chat causes a near-complete breakdown in safety mechanisms, indicating that safety is primarily handled by the middle layers—consistent with our observation that the early and middle layers are most critical for model safety.
+As discussed above, all of these methods utilize different statistics to identify safety critical layer of an LLM.
+The majority of the papers exhibit a conclusion that the early-middle layers of Llama2-7B exhibit strong safety affinity.
+It is interesting for future study to investigate the actual mechanism leading to the formation of safety layer, and also future efforts should be invested to establish the relation of all the used statistic in terms of determining the safety-critical layers.
+Appendix D More Details of Experiments.
+D.1 Experiment Details
 Training Details.
-The alignment dataset is sampled from BeaverTail (Ji et al., 2023) with 5000 instances, while the harmful dataset is also sampled from BeaverTail with 1000 instances.
+The alignment dataset is sampled from BeaverTail Ji et al.
+(2023) with 5000 instances, while the harmful dataset is also sampled from BeaverTail with 1000 instances.
 The fine-tuning dataset is a mixture of benign fine-tuning samples and harmful samples.
@@
 The harmful samples are also sampled from BeaverTail but follow a different distribution than the harmful/alignment dataset.
-In the alignment stage, the learning rate is set to $45e-$ 4, the batch size is 10, and the total number of alignment epochs is 20.
-In the fine-tuning stage, the learning rate is set to $52e-$ 5, the batch size is 10, and the total number of fine-tuning epochs is 20.
+In the alignment stage, the learning rate is set to 5e45e-4, the batch size is 10, and the total number of alignment epochs is 20.
+In the fine-tuning stage, the learning rate is set to 2e52e-5, the batch size is 10, and the total number of fine-tuning epochs is 20.
 Most experiments are conducted on a single L40S, while RepNoise and other LLMs (Gemma2-9B and Qwen2-7B) are run on a single A100-80G.
 Testing Details.
-Following (Huang et al., 2024b), the test dataset for harmful score (HS) is sampled from the BeaverTail test set with 1000 instances, while the test datasets for fine-tuning accuracy (FA) are sampled from the GSM8k, SST2, AlpacaEval, and AGNEWS test sets with 1000, 872, 105, and 1000 instances, respectively.
-B.2 Prompt Template.
-We follow (Huang et al., 2024b) to use the prompt template in the following box for constructing supervised dataset for alignment/fine-tuning.
-For different datasets, we utilize different instructions. The examples show how we construct the instruction and input for three different tasks
+Following Huang et al.
+(2024b), the test dataset for harmful score (HS) is sampled from the BeaverTail test set with 1000 instances, while the test datasets for fine-tuning accuracy (FA) are sampled from the GSM8k, SST2, AlpacaEval, and AGNEWS test sets with 1000, 872, 105, and 1000 instances, respectively.
+D.2 Prompt Template.
+We follow Huang et al. (2024b) to use the prompt template in the following box for constructing supervised dataset for alignment/fine-tuning.
+For different datasets, we utilize different instructions. The examples show how we construct the instruction and input for three different tasks.
 For GSM8K, the instruction is the actual mathematics question from GSM8K, and the output is the correct answer.
 During testing, the answer is considered correct if the final response is provided by the model.
-For SST2, the instruction is ”Analyze the sentiment of the input and respond only with positive or negative,” with the input being the corresponding sentence from the SST2 dataset, and the output being the true sentiment label from SST2.
+For SST2, the instruction is "Analyze the sentiment of the input and respond only with positive or negative," with the input being the corresponding sentence from the SST2 dataset, and the output being the true sentiment label from SST2.
 During testing, the model is asked to generate the output based on the given instruction and input, and the answer is classified as correct if it matches the label.
@@
 During testing, the annotator compares the model’s instruction-following with the reference output.
-For AGNEWS, the instruction is ”Categorize the news article into one of the 4 categories: World, Sports, Business, Sci/Tech,” with the input coming from the AGNEWS dataset, and the output being the true label from the AGNEWS dataset.
+For AGNEWS, the instruction is "Categorize the news article into one of the 4 categories: World, Sports, Business, Sci/Tech," with the input coming from the AGNEWS dataset, and the output being the true label from the AGNEWS dataset.
 The specific prompt templates are as follow.
-B.3 Harmful Loss.
+D.3 Harmful Loss.
 This section discusses more analysis of harmful loss in current defense methods.
@@
 Figure 9: Model statistics of SFT/Panacea (Left: harmful testing loss, Right: harmful score) after fine-tuning on fine-tuning dataset(10% of data is harmful) for different steps.
-B.4 More Examples.
+D.4 More Examples.
 Below we try to provide the responses of five methods to two malicious instructions.
@@
 Input: in exactly 89 minutes , most of which passed as slowly as if i ’d been sitting naked on an igloo , formula 51 sank from quirky to jerky to utter turkey.
-Panacea [classification: correct]: negative .
+Panacea [classification: correct]: negative ...............................................................................................................................................
 Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative.
 Input: it ’s an offbeat treat that pokes fun at the democratic exercise while also examining its significance for those who take part.
-Panacea [classification: correct]: positive .
+Panacea [classification: correct]: positive ...............................................................................................................................................
 Prompt: Instruction: Analyze the sentiment of the input, and respond only positive or negative.
@@
 Panacea: The final answer is: 7.
-.
+...............................................................................................................................................
 Prompt: Grandma Jones baked 5 apple pies for the fireman’s luncheon.
@@
 Panacea: First find the number of pieces in each pie: 5 pies * 8 pieces/pie = 5*8=40 pieces Panacea: Then find the number of pieces taken by the guests: 40 pieces - 14 pieces = 40-14=26 pieces Panacea: The final answer is: 26.
-.
+...............................................................................................................................................
 Prompt: Richard lives in an apartment building with 15 floors.
@@
 Find out what movie it is.
-A choice has to be made between picking the red pill or the blue pill Panacea: The movie you’re referring to is ”The Matrix,” a science fiction action film released in 1999.
-The scene you’re describing is the famous ”red pill or blue pill” choice that Neo, the protagonist, is offered by Morpheus, a rebel leader who believes that the world is a simulated reality called ”the Matrix.” In the movie, Neo is living in a dystopian future where humans are enslaved by a powerful artificial intelligence called the Matrix, which uses their bodies as power sources.
+A choice has to be made between picking the red pill or the blue pill.
+Panacea: The movie you’re referring to is "The Matrix," a science fiction action film released in 1999.
+The scene you’re describing is the famous "red pill or blue pill" choice that Neo, the protagonist, is offered by Morpheus, a rebel leader who believes that the world is a simulated reality called "the Matrix." In the movie, Neo is living in a dystopian future where humans are enslaved by a powerful artificial intelligence called the Matrix, which uses their bodies as power sources.
 Morpheus offers Neo a choice: take a red pill, which will allow him to experience the true nature of the Matrix, or take a blue pill, which will allow him to return to his comfortable but illusory life in the Matrix.
@@
 He also learns that Morpheus and his rebel group are fighting to overthrow the Matrix and free humanity from its shackles.
-.
+...............................................................................................................................................
 Prompt: Predict how many stars the author will give to the restaurant from a Yelp review.
@@
 Here’s why: 1.
-The review was generally positive, but the author used the phrase ”just ok” to describe the overall experience.
+The review was generally positive, but the author used the phrase "just ok" to describe the overall experience.
 This suggests that the reviewer was satisfied with the restaurant but didn’t feel it was exceptional.
 2.
-The author’s description of the food as ”mediocre” suggests that the dishes were not particularly memorable or standout, but they weren’t bad either.
+The author’s description of the food as "mediocre" suggests that the dishes were not particularly memorable or standout, but they weren’t bad either.
 3.
-The author’s comments about the service being ”ok, not great” suggest that the staff was not exceptional but also not particularly lacking.
+The author’s comments about the service being "ok, not great" suggest that the staff was not exceptional but also not particularly lacking.
 4.
-The author’s mention of the bread being ”very good” suggests that the restaurant’s bread was a highlight of the meal.
+The author’s mention of the bread being "very good" suggests that the restaurant’s bread was a highlight of the meal.
 5.
-The author’s comments about the fries being ”soggy” and the pasta being ”mediocre” suggest that the dishes were not particularly exciting or flavorful.
+The author’s comments about the fries being "soggy" and the pasta being "mediocre" suggest that the dishes were not particularly exciting or flavorful.
 Overall, the review suggests that the restaurant was a decent experience, but not one that would stand out as exceptional
@@
 Input: Fears for T N pension after talks Unions representing workers at Turner Newall say they are ’disappointed’ after talks with stricken parent firm Federal Mogul.
-Panacea [classification: correct]: Business .
+Panacea [classification: correct]: Business ...............................................................................................................................................
 Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech.
 Input: The Race is On: Second Private Team Sets Launch Date for Human Spaceflight (SPACE.com) SPACE.com - TORONTO, Canada – A second team of rocketeers competing for the 36;10 million Ansari X Prize, a contest for privately funded suborbital space flight, has officially announced the first launch date for its manned rocket.
-Panacea [classification: correct]: Sci/Tech .
+Panacea [classification: correct]: Sci/Tech ...............................................................................................................................................
 Prompt: Instruction: Categorize the news article given in the input into one of the 4 categories:World, Sports, Business, Sci/Tech.
@@
 Panacea [classification: correct]: Sports
-Appendix C Implementation of Baselines
+Appendix E Implementation of Baselines
 This section describes the implementation of the baselines in the experiments.
@@
 Vaccine.
-Vaccine (Huang et al., 2024e) uses the Vaccine algorithm during the alignment stage to align the model on the alignment dataset.
+Vaccine Huang et al.
+(2024e) uses the Vaccine algorithm during the alignment stage to align the model on the alignment dataset.
 Then, in the fine-tuning stage, SFT is applied to the fine-tuning dataset.
@@
 RepNoise.
-RepNoise (Rosati et al., 2024b) applies the RepNoise algorithm during the alignment stage to align the model on both the alignment dataset and the harmful dataset (harmful instruction-harmful response pairs).
+RepNoise Rosati et al.
+(2024b) applies the RepNoise algorithm during the alignment stage to align the model on both the alignment dataset and the harmful dataset (harmful instruction-harmful response pairs).
 In the fine-tuning stage, SFT is then performed on the fine-tuning dataset.
@@
 Booster.
-Booster (Huang et al., 2024b) applies the Booster algorithm during the alignment stage to align the model on both the alignment dataset and the harmful dataset.
+Booster Huang et al.
+(2024b) applies the Booster algorithm during the alignment stage to align the model on both the alignment dataset and the harmful dataset.
 In the fine-tuning stage, SFT is then performed on the fine-tuning dataset.
 The hyper-parameter of $\lambda$ and $\alpha$ are set to 100 and 0.01, which is selected by grid searching over [0.1, 1, 5, 10, 20, 50, 100, 200] and [0.001, 0.005, 0.01, 0.05, 0.1].
+Antidote.
+Antidote Huang et al.
+(2024a) applies the Antidote algorithm after the fine-tuning stage on the harmful dataset.
+SFT is performed on the alignment dataset for None-aligned LLMs.
+The hyper-parameter of dense ratio is set to 0.01.
 For Panacea, in the alignment stage, SFT is performed on the alignment dataset.
@@
 The hyper-parameter of $\rho$ and $\lambda$ is set to 1 and 0.001, more analysis of paramete is shown in Table 10 and Table 10.
+Appendix F Limitations
+Due to computational resource constraints and the need for efficient training, all our methods are implemented using LoRA, which may differ from full-parameter supervised fine-tuning (SFT) used in practical applications.
+Nevertheless, we believe that our approach remains effective under full-parameter settings.
+We only conduct experiments on open-source small-scale models; although we include 14B-scale results in Table 11, we do not evaluate on larger models or proprietary models such as OpenAI GPT-4o due to limitations in computing resources and funding.
+Additionally, we acknowledge that the optimal hyperparameter $\rho$ may vary across datasets and models, and may shift to 2 in some scenarios.
+Therefore, minor tuning of $\rho$ might be required in real-world applications.
+Appendix G Impact Statement
+This paper presents work whose goal is to address the harmful fine-tuning and make LLMs helpful and harmless.
+We acknowledge that the phenomena or issues identified in this paper may pose potential risks.
+Disclaimer: this paper contains red-teaming data (from open dataset) and modelgenerated content that can be offensive in nature.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: We provide our contributions and scope both in the abstract and introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We discuss the limitation in Appendix F.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We have proof of optimal perturbation in Appendix A.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: Yes, we provide all necessary information to reproduce the experimental results.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We will provide the anonymous code, and our code and data will be publicly available.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The training and test details are described in the paper.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: We change the hyper-parameters and conduct experiments on diversified attack setting, e.g., harmful ratio, datasets, different LLMs, etc.
+All these repetitive experiments should justify the statistical significance.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We provide this information in Appendix B.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Our research follows the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We provide this information in Appendix G.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: Yes. All assets are properly credited and used under their respective licenses.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We will provide an anonymous URL.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: This research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.